### v1

# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

Abstract
Gurumuhkani et al. (CCC’24) introduced the local enumeration problem Enum($k$, $t$) as follows: for a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF with no satisfying assignment with Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$. They showed that efficient algorithms for local enumeration yield new $k$-SAT algorithms and depth-3 lower bounds for Majority function. As the first non-trivial case, they gave an algorithm for $k=3$ which in particular gave a new lower bound on the size of depth-3 circuits with bottom fan-in at most 3 computing Majority. In this paper, we give an optimal algorithm that solves local enumeration on monotone formulas for $k=3$ and all $t\leq n/2$. In particular, we obtain an optimal lower bound on the size of monotone depth-3 circuits with bottom fan-in at most 3 computing Majority.
1 Introduction
Ajtai [1] and Furst, Saxe, and Sipser [2] showed that the Parity function, which is easily computable in linear time, cannot be computed by constant-depth circuits of polynomial size. Later, Håstad [8] proved his landmark Switching Lemma to show a tight exponential lower bound: Parity requires depth-$d$ circuits of size $2^{\Omega(n^{{1}/{(d-1)}})}$. Four decades later, this result remains the state-of-the-art for ${\mathbf{AC}}^{0}$ lower bounds. Even for the seemingly simple case of depth-3 circuits, we do not have a lower bound of $2^{\omega(\sqrt{n})}$ for an explicit function. This bound has become a long-standing barrier, and surpassing it is recognized as a major open problem in the field.
Depth-3 circuits are interesting in their own right. Valiant [17] showed that any linear-size circuit of logarithmic depth can be computed by a depth-3 circuit of size $2^{O(n/\log\log n)}$. Thus, strong enough depth-3 lower bounds imply super-linear lower bounds for circuits of logarithmic depth. A particularly interesting class of depth-3 circuits is $\Sigma_{3}^{k}$; Or-And-Or circuits where the bottom gates have a fan-in of at most $k$, i.e., disjunctions of $k$-CNFs. As a natural measure of size, we denote the minimum top fan-in of a $\Sigma^{k}_{3}$ circuit computing $f$ by $\operatorname{Size}^{k}_{3}(f)$, and by $\operatorname{Size}^{+k}_{3}(f)$ when the circuit is monotone. Golovnev, Kulikov, and Williams [3] recently revisited Valiant’s approach and showed that even this restricted class of depth-3 circuits capture general circuits; a near-maximal $2^{n-o(n)}$ lower bound for $\Sigma^{16}_{3}$ implies new lower bounds for unrestricted circuits.
$\Sigma^{k}_{3}$ lower bounds.
Paturi, Pudlák, and Zane [13] (henceforth called PPZ) showed that Parity requires $\Sigma^{k}_{3}$ circuits of size $\Omega(2^{n/k})$. In fact, this results in an unrestricted depth-3 lower bound of $\Omega(n^{1/4}2^{\sqrt{n}})$. Both of these lower bounds are known to be tight. Paturi, Pudlák, Saks, and Zane [12] (henceforth called PPSZ) extended the techniques of [13] and showed that recognizing the codewords of a good error-correcting code such as BCH requires $\Sigma^{k}_{3}$ circuits of size $2^{cn/k}$ for a universal constant $c>1$. This remains the best known lower bound for computing any explicit function by $\Sigma^{k}_{3}$ circuits. The core technique in these lower bounds is a combinatorial result about the structure of satisfying assignments of $k$-CNF formulas which can be exploited to solve $k$-SAT.
Non-trivial $k$-SAT algorithms.
An algorithm solving $k$-SAT is non-trivial if it has running time $2^{(1-\epsilon_{k})n}$ for some $\epsilon>0$ which we call savings. PPZ and PPSZ used their combinatorial arguments to give non-trivial $k$-SAT algorithms. PPSZ currently holds the record as the best known $k$-SAT algorithm with savings $c/k$ for a universal constant $c>1$ [15]. The Super Strong Exponential Time Hypothesis (SSETH) states that such savings, that is $\Theta(1/k)$, are best possible [18]. We know that indeed this is the case for PPSZ [14].
Depth-3 complexity of Majority function.
The Majority function has long been held as a natural candidate to beat the depth-3 barrier [7]. Unlike Parity, the tightness of natural circuit constructions for Majority is not known. To determine the depth-3 complexity of Majority, Gurumukhani et al. [5] suggested a path forward by considering $\Sigma^{k}_{3}$ circuits. Their central conjecture is that the size of any $\Sigma_{3}^{k}$ circuit computing Majority is $2^{\Omega(n\log(k)/k)}$. Together with a standard random restriction argument to reduce the bottom fan-in, this conjecture implies an unrestricted depth-3 lower bound of $2^{\Omega(\sqrt{n\log n})}$ matching a known upper bound [10]. They pointed out that determining $\Sigma^{k}_{3}$ complexity of Majority is non-trivial even for $k=3$, and proposed this as a first step towards settling the depth-3 complexity of Majority. Results of Lecomte, Ramakrishnan, and Tan [11] can be interpreted as proving the conjecture under the assumption that each depth-2 subcircuits depends on at most $k$ input bits.
Similar to PPZ and PPSZ, determining the depth-3 complexity is closely tied to $k$-SAT. Gurumukhani et al. [6] showed that proving the conjectured lower bound for $\Sigma^{k}_{3}$ via a particular approach which they called local enumeration, would break SSETH.
Local enumeration.
Enum($k$, $t$) is defined as follows. For a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF $F$ with no satisfying assignment of Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$. Analogously, we can define a restriction of this problem to monotone formulas which we denote by Enum+($k$, $t$).
The connection between Majority, $k$-SAT, and local enumeration is given by the following theorem.
Theorem 1.1.
[5] Assume that Enum($k$, $t$) can be solved in (expected) time $b(n,k,t)$. Then
  1. 1.
    $\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}})\geq\binom{n}{n/2}/b(n,k,n/2)$,
  2. 2.
    $k$-SAT can be solved in (expected) time $O(\sum_{t=0}^{n/2}b(n,k,t))$.
Gurumukhani et al. [5] considered $k=3$ as the first non-trivial instance and showed that Enum($3$, $n/2$) can be solved in expected time $1.598^{n}$. Gurumukhani et al. [6] later showed that solving a restriction of Enum($k$, $t$) in which we must enumerate only the Not-All-Equal (NAE) solutions is enough to beat the depth-3 barrier and SSETH. They gave an algorithm that optimally solves this problem for $k=3$.
1.1 Our contribution
Our main result is an optimal lower bound on the size of any monotone $\Sigma^{3}_{3}$ circuit computing Majority matching the upper bounds conjectured in the literature (see e.g. [4]).
Theorem 1.2.
$\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq 2^{O(n\log(k)/k)}$. In particular, $\operatorname{Size}^{3}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq(2/6^{1/4})^{n}\simeq 1.277^{n}$. Furthermore, there are monotone circuits that achieve this bound.
We optimally solve Enum+($3$, $n/2$) to derive our lower bound
Theorem 1.3 (Main result).
There exists a deterministic algorithm that solves Enum+($3$, $n/2$) in time ${\mathtt{poly}}(n)\cdot 6^{n/4}$. Consequently, $\operatorname{Size}^{+3}_{3}(\operatorname{\mathsf{Maj}}_{n})\geq(2/6^{1/4})^{n-o(n)}\simeq 1.277^{n}$.
While we do not claim that our techniques are new, we believe that our approach to prove matching lower bounds for depth-3 circuits holds promise. The first element of our approach is the top-down technique: if a $\Sigma\Pi\Sigma_{3}^{k}$ circuit computes a Boolean function $f:\{0,1\}^{n}\mapsto\{0,1\}$ and $s$ is the number of its depth-2 circuits, one of the depth-2 circuits which is a $k$-CNF must accept at least $|f^{-1}(1)|/s$ members of $f^{-1}(1)$. This brings our focus to the study of extremal $k$-CNFs with respect to $f$. A $k$-CNF $F$ is extremal with respect to $f$ if ${\textsc{Sat}}(F)\subseteq f^{-1}(1)$ and $|{\textsc{Sat}}(F)|$ is the maximum among such $k$-CNF. Such extremal $k$-CNFs can also used for constructing depth-3 $\Sigma\Pi\Sigma_{3}^{k}$ circuits which are (almost) optimal depending on the symmetries of $f$. Thus, the study of $k$-CNF which are extremal with respect to a Boolean function is of fundamental concern.
What do we know of $k$-CNFs which are extremal with respect to the majority function or more generally with respect to Boolean threshold functions? Our intuition (based on extensive provably optimal constructions for small $n$) suggested that a modularity principle might hold. We say that modularity principle holds for a Boolean function if an extremal $k$-CNF $F$ can be expressed as the conjunction of $k$-CNFs $F^{\prime}_{i}$ where $F^{\prime}_{i}$ is supported on a bounded number of variables and the supports of $F^{\prime}_{i}$ are disjoint. Although we lack a formal method to determine if the majority function is inherently modular, we used the modularity principle to guide a trial-and-error search for constructing putative extremal 3-CNFs of Boolean threshold functions (for thresholds less or equal to $n/2$).
Our final step is to apply induction to prove the optimality of our constructions. Several ingredients helped the induction to succeed: 1. Correct guess (or construction) of optimal constructions for Boolean threshold function with threshold less or equal to $n/2$, 2. Construction of (optimal or near optimal) extremal $3$-CNFs which are further restricted in certain ways (for example, $3$-CNFs which are required to include a $2$-clause), and 3. identification of appropriate anchors to carry out the induction taking advantage of forced variables (this ingredient in explain in detail in Section 4. In a nutshell, starting with precisely optimal constructions played an essential role in the proving optimality of our constructions.
2 Set Systems
A set system $\mathbb{F}=(V,F)$ is a collection $F$ of sets over a set $V$ of elements. If $V$ is a set of vertices, then $\mathbb{F}$ is interpreted as a graph. If $V$ is as a set of Boolean variables, then $\mathbb{F}$ is interpreted as a monotone CNF. While we regard $\mathbb{F}$ as a set system most of the time, we invite the reader to apply other interpretations to gain additional viewpoints. When we refer to sets in $\mathbb{F}$, we mean the sets in $F$.
Notation 2.1.
We use $n(\mathbb{F})$ to denote the number of elements and $m(\mathbb{F})$ to denote the number of sets of $\mathbb{F}$. We use $k$-set to refer to a set with cardinality $k$. We extend the notation to $k$-clauses and $k$-edges.
We often work with set systems that where the size of the sets is bounded. We say that a set system $\mathbb{F}=(V,F)$ is a $k$-system if each $C\in F$ has cardinality at most $k$. If every set in the set system is a $k$-set, then we call it a $k$-uniform set system.
We now introduce the essential concept of a transversal of a set system.
Definition 2.2 (Transversals).
Let $\mathbb{F}=(V,F)$ be a set system. An $\ell$-transversal of $\mathbb{F}$ is a subset of $V$ of size $\ell$ that intersects every member of $F$. We use transversal to refer to an $\ell$-transversal for some $\ell\geq 0$.
Fact 2.3 (Critical sets of a transversal [13]).
If $\gamma$ is a minimum-size transversal of a set system $\mathbb{F}$, for every $x\in\gamma$ there exists a set $S_{x}$ in $\mathbb{F}$ such that $S_{x}\cap\gamma=\{x\}$. We call $S_{x}$ a critical set for $x$ with respect to $\gamma$.
We parameterize set systems by the minimum size of its transversals to arrive at the notion of $t$-threshold systems.
Definition 2.4 ($t$-threshold set systems).
A set system is called $t$-threshold if all its transversals have size at least $t$.
We occasionally need the following concept.
Definition 2.5 (Transversal number).
The transversal number $\tau(\mathbb{H})$ of a set system $\mathbb{H}$ is the cardinality of its minimum-size transversals.
Notation 2.6.
We use $\gamma_{t}(\mathbb{F})$ to denote the set of $t$-transversals of $\mathbb{F}$ and $\theta_{t}(\mathbb{F})$ to denote the number of $t$-transversals of $\mathbb{F}$.
Definition 2.7 (Extremal $t$-threshold $k$-set systems).
We say that a $t$-threshold $k$-system $\mathbb{F}$ is extremal if $\theta_{t}(\mathbb{F})$ is the maximum over all $t$-threshold $k$-systems on $n(\mathbb{F})$ elements.
Definition 2.8.
For $n,k\geq 1$ and $0\leq t\leq n$, we define $\Theta(n,t,k)$ to be equal to $\theta_{t}(\mathbb{H})$ where $\mathbb{H}$ is an extremal $t$-threshold $k$-system over $n$ elements.
The following states that uniform set systems suffice as far as extremal $t$-threshold $k$-set systems are concerned.
Fact 2.9.
For $t\leq n-k+1$ and for every $t$-threshold $k$-system, there is a $t$-threshold $k$-uniform set system which has the same set of $t$-transversals. For t=nk+it=n-k+i and $1\leq i\leq k$, the $t$-threshold (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is extremal.
2.1 Circuits and Transversals
$k$-systems play a key role in the (monotone) circuit complexity of Boolean functions. Our circuits are $\Sigma\Pi\Sigma_{k}$ circuits which are depth-3 circuits where the top gates is an $\mathbf{OR}$ gate and its depth-2 subcircuits are $k$-CNFs. The size of such a circuit is the number of its depth-2 subcircuits. For monotone $\Sigma\Pi\Sigma_{k}$ circuits, the depth-2 subcircuits are $k$-systems.
Definition 2.10.
A Boolean function $f$ on $n$ variables $x_{1},\ldots,x_{n}$ is a $t$-threshold function for $0\leq t\leq n$ provided $f(x_{1},\ldots,x_{n})=1$ if and only if $\Sigma_{i}x_{i}\geq t$.
If a $k$-system $\mathbb{F}$ is a depth-2 subcircuit of a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function, then $\mathbb{F}$ is a $t$-threshold system. Since no such $\mathbb{F}$ has more than $\Theta(n,t,k)$ $t$-transversals, we conclude the following.
Fact 2.11.
If a monotone $\Sigma\Pi\Sigma_{k}$ circuit computes a $t$-threshold function, then its size is at least $\binom{n}{t}/\Theta(n,t,k)$.
Furthermore, it is known that we only need extremal $t$-threshold $k$-systems as depth-2 subcircuits to construct size-optimal (up to a polynomial factor) monotone $\Sigma\Pi\Sigma_{k}$ circuits for computing $t$-threshold functions. For every $t$-threshold $k$-system $\mathbb{F}$, it is known that there exists a $\Sigma\Pi\Sigma_{k}$ circuit of size ${\mathtt{poly}}(n)\binom{n}{t}/\theta(\mathbb{F})$ for computing a $t$-threshold function. This can be shown by exploiting the symmetries of the threshold functions and monotonicity [4].
Fact 2.12.
There exists a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function whose size is at most ${\mathtt{poly}}(n)\binom{n}{t}/\Theta(n,t,k)$.
2.2 Connection to Turán Systems
The following definition is adapted from [16, 9].
Definition 2.13 (Turán system and Turán number).
For integers $n\geq\ell\geq k\geq 1$, a $k$-system over an $n$-element set is called an $(n,\ell,k)$-Turán system if every $\ell$-set contains a set of the system. The Turán number ${\Psi}(n,\ell,k)$ is defined as
${\Psi}(n,\ell,k)=\max_{j\leq k}\min m(\mathbb{F}),$
where the minimum is taken over all $(n,\ell,j)$-Turán systems $\mathbb{F}$.
Fact 2.14.
Assume that $2k\leq n-1$. A collection of $k$-sets of an $n$-element set is an $(n,k+1,k)$-Turán system if and only if it is a (nk)(n-k)-threshold $k$-set system.
Proof.
Let $\mathbb{F}$ be an $(n,k+1,k)$-Turán system. Since every $(k+1)$-set includes a $k$-set of $\mathbb{F}$, every (nk1)(n-k-1)-set fails to intersect some member of $\mathbb{F}$. Thus $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-uniform set system.
Conversely, if $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-set system, then no (nk1)(n-k-1)-set is a transversal for $\mathbb{H}$. By considering complements, we conclude $\mathbb{F}$ must then be an $(n,k+1,k)$-Turán system. ∎
Furthermore, we have
Fact 2.15.
$\Theta(n,n-k,k)\;=\;\binom{n}{k}\;-\;{\Psi}(n,k+1,k).$
2.3 Configurations
To prove upper bounds on the number of $t$-transversals of a $t$-threshold system, our arguments will be based on whether certain configurations of sets are present in the set system. In the following, we catalog these configurations.
Definition 2.16 (Triangle configuration).
We say that a set system contains a triangle configuration if it contains three sets of the form $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
Definition 2.17 ( $\mathbf{\exists}$} configuration).
We say that a set system contains a $\mathbf{\exists}$} configuration if it contains four sets of the form $\{a,b,c\}$, $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ where the set $\{d,e,f,g,h,i\}$ is disjoint from the set $a,b,c$. We say that $\{a,b,c\}$ is the spine and $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ are the arms of the $\mathbf{\exists}$} configuration.
We sometimes regard a collection of $2$-sets as a graph where each $2$-set is an edge and use graph-theoretic terminology to describe configurations.
Definition 2.18 (Path configuration).
We say that a collections of $2$-sets is a path of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{\ell}\}$ where $a_{0},a_{1},\ldots,a_{\ell}$ are distinct.
Definition 2.19 (Cycle configuration).
We say that a collection of $2$-sets is a cycle of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{0}\}$ where $a_{0},a_{1},\ldots,a_{\ell-1}$ are distinct.
3 Constructions for Extremal Threshold System
In this section, we construct $k$-systems over $n\geq 1$ elements for transversal number $0\leq t\leq\lfloor\frac{n}{2}\rfloor$. In the next section, we argue that they are $t$-extremal. We first consider certain constructions for bounded $n$ which will serve as building blocks for general constructions. The basic building blocks come in two flavors: cliques and Turán-style constructions.
3.1 Cliques
Definition 3.1 (Clique).
For $l\geq k\geq 1$, let ${\mathbb{K}}_{l}^{k}$ denote the $k$-system of all $k$-sets on $l$ elements.
The following two cliques are widely used in our constructions: ${\mathbb{K}}_{3}^{3}=\{\{1,2,3\}\}$ is a $1$-threshold 33-system where $\theta_{1}({\mathbb{K}}_{3}^{3})=3$; ${\mathbb{K}}_{4}^{3}=\{\{1,2,3\},\{1,2,4\},\{1,3,4\},\{2,3,4\}\}$ is a $2$-threshold 33-system where $\theta_{2}({\mathbb{K}}_{4}^{3})=6$.
Fact 3.2.
${\mathbb{K}}_{l}^{k}$ is an extremal (lk+1)(l-k+1)-threshold $k$-uniform set system on $l$ elements where $n({\mathbb{K}}_{l}^{k})=l,$ $m({\mathbb{K}}_{l}^{k})=\binom{l}{k}$, $\tau({\mathbb{K}}_{l}^{k})=l-k+1,$ and $\Theta(l,l-k+1,k)=\theta_{l-k+1}({\mathbb{K}}_{l}^{k})=\binom{l}{l-k+1}$.
3.2 Turán-style constructions
Let $n\geq 3$. Constructions from here on are primarily limited to $k=3$. Consider the following (n,n3,3)(n,n-3,3)-Turán system ${\mathbb{T}}_{n}^{3}$ which is obtained by what we call a TuránTur\'{a}n-style construction [16]. Partition $n$ into three parts where the parts have sizes as equal as possible and organize them in a circular fashion. ${\mathbb{T}}_{n}^{3}$ consists of the following $3$-sets:
  1. 1.
    For each part, all 3-sets of the part
  2. 2.
    For each part, all 3-sets of the form where two elements come from the part and one element from the adjacent part in the circular order.
Conjecture 3.3 (Turán).
[…]
The conjecture implies
Fact 3.4.
Conjecture 3.3 implies
[…]
We use the following Turán-style building blocks in our constructions.
Example 3.5.
${\mathbb{T}}_{5}^{3}=\{\{1,2,3\},\{3,4,5\},\{1,2,4\}\}$ is a $2$-threshold 33-uniform set system, $n({\mathbb{T}}_{5}^{3})=5$, $m({\mathbb{T}}_{5}^{3})=3$, $\tau({\mathbb{T}}_{5}^{3})=2$, and $\theta_{2}({\mathbb{T}}_{5}^{3})=7$.
Example 3.6.
${\mathbb{T}}_{6}^{3}=\{\{x_{0},y_{0},x_{1}\},\{x_{0},y_{0},y_{1}\},\{x_{1},y_{1},x_{2}\},\{x_{1},y_{1},y_{2}\},\{x_{2},y_{2},x_{0}\},\{x_{2},y_{2},y_{0}\}\}$ is a $3$-threshold 33-uniform set system, $n({\mathbb{T}}_{6}^{3})=6$, $m({\mathbb{T}}_{6}^{3})=6$, $\tau({\mathbb{T}}_{6}^{3})=3$, and $\theta_{3}({\mathbb{T}}_{6}^{3})=14$.
Proofs of the following statement are provided in the appendix (Theorems A.1 and A.2).
Fact 3.7.
$\Theta(5,2,3)$ $=\theta_{2}({\mathbb{T}}_{5}^{3})=7$
$\Theta(6,3,3)$ $=\theta_{3}({\mathbb{T}}_{6}^{3})=14$
3.3 Constructions for $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$
In this section, we provide constructions that work for all $n,k\geq 1$, $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$. We need the following notation to represent larger set systems composed of smaller set systems. If $\mathbb{F}_{1}=(V_{1},F_{1})$ and $\mathbb{F}_{2}=(V_{2},F_{2})$ are $k$-systems over disjoint sets of elements, the sum $\mathbb{F}_{1}+\mathbb{F}_{2}$ of $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ represents the $k$-system $\mathbb{F}=(V_{1}\cup V_{2},F_{1}\cup F_{2})$. For a $k$-system $\mathbb{F}$, $\ell\cdot\mathbb{F}$ represents the $k$-system obtained by summing $\ell$ disjoint copies of $\mathbb{F}$.
Let $\mathbb{S}^{k}_{n,t}=t\cdot{\mathbb{K}}_{k}^{k}$. For $n\geq 1$ and $0\leq t\leq\lfloor n/k\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique optimal construction for any $k\geq 1$.
Theorem 3.8 ([4]).
For $k,n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{k}\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique (up to isomorphism) extremal $t$-threshold $k$-system on $n$ elements.
For t=nk+it=n-k+i where $1\leq i\leq k$, we have
Fact 3.9.
The (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is an extremal $t$-threshold $k$-system and $\Theta(n,n-k+i,k)=\binom{n}{n-k+i}$.
3.4 Constructions for $n\geq 1,k=3$ and $0\leq t\leq\lfloor\frac{n}{2}\rfloor$
Our constructions for $n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{2}\rceil$ are limited to the case $k=3$. They use ${\mathbb{K}}_{3}^{3}$, ${\mathbb{K}}_{4}^{3}$, ${\mathbb{T}}_{5}^{3}$, and ${\mathbb{T}}_{6}^{3}$ as building blocks. The following important reparameterization is useful to express our constructions.
Definition 3.10.
For $k=3$, we rewrite n=3tsn=3t-s where $s$ is called the deficit of the elements (variables or vertices) with respect to the threshold $t$. To emphasize the role of $s$, we specialize $\Theta(n,t,k)$ for $k=3$ and use the function $\Phi_{0}(s,t)$ to denote the maximum number of $t$-transversals of a $t$-threshold 33-system with deficit $s$. In other words, $\Phi_{0}(s,t)=\Theta(3t-s,t,3)$.
Let $s\leq t$ and $t\geq 0$ (equivalently $n\geq 2t\geq 0$). Let $\mathbb{P}^{3}_{s,t}$ be the $3$-system on n=3tsn=3t-s elements defined as follows.
[…]
Fact 3.11.
$\mathbb{P}^{3}_{s,t}$ is a $t$-threshold 33-uniform set system.
Fact 3.12.
[…]
Remark 3.13.
${\mathbb{K}}_{3}^{3}+{\mathbb{T}}_{6}^{3}$ can be replaced ${\mathbb{K}}_{4}^{3}+{\mathbb{T}}_{5}^{3}$ with out affecting the number of minimum-size transversals.
We summarize the results of this section by the following theorem which provides lower bounds for the maximum number of minimum-size transversals $\Phi(s,t)$ of $3$-systems over n=3tsn=3t-s elements with transversal number and deficit satisfying the conditions, $s\leq t$ and $t\geq 0$.
Theorem 3.14.
[…]
Proof.
The first equation follows from [4]. The second equation is straightforward. The remaining two inequalities follow from Fact 3.12. ∎
4 Upper Bounds on $\Phi_{0}(s,t)$
In this section, we establish matching upper bounds on the number of transversals $\Phi_{0}(s,t)$ of extremal $t$-threshold 33-systems for $t\geq 0$ and $s\leq t$. We prove these bounds by induction on 3ts3t-s and $t$. As we do induction on $s$ and $t$, we have to deal with restricted set systems which motivates the following definitions.
For $i\in\{1,2,3,4\}$, we say a $3$-system $\mathbb{F}$ is a type $i$ 33-system if it contains at least $i$ 22-sets. We further specialize the type 2 systems: $\mathbb{F}$ is a type $2o$ 33-system if it contains two $2$-sets which have exactly one element in common and a type $2d$ 33-system if it contains two $2$-sets which are disjoint. We also restrict type 3 systems as follows: $\mathbb{F}$ is a type $3x$ 33-system if it contains three $2$-sets which do not have an element in common. By extension, unrestricted $3$-systems are of type 0. For $i\in\{1,2d,2o,3x,4\}$, let
  • ${\Xi}_{i}$ denote the family of $3$-system of type $i$,
  • ${\Xi}_{i}(s,t)$ denote the family of $t$-threshold 33-system of type $i$ over n=3tsn=3t-s elements, and
  • $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold 33-systems over 3ts3t-s elements.
We prove upper bounds on $\Phi_{0}(s,t)$ by simultaneously proving upper bounds on $\Phi_{i}(s,t)$ by induction on 3ts3t-s and $t$. Our theorem on upper bounds is presented below.
Theorem 4.1.
The following upper bounds hold for $\Phi_{i}(s,t)$ for $i\in\{1,2d,2o,3,4\}$.
3-systems
[…]
3-systems with at least one $2$-set
[…]
3-systems with two $2$-sets which overlap in one element
[…]
3-systems with two $2$-sets which are disjoint
[…]
3-systems with three $2$-sets which do not have an element in common
[…]
3-systems with four $2$-sets
[…]
4.1 Outline of the proof
The proof assumes that $t\geq 1$ as the theorem is trivially true otherwise. We divide the task of proving the upper bounds into six parts. In part $i\in\{0,1,2d,2o,3x,4\}$, we prove the upper bounds for $\Phi_{i}(s,t)$, the number of $t$-transversals for set systems in ${\Xi}_{i}(s,t)$ assuming the upper bounds for $\Phi_{j}(s,t)$ for smaller values of 3ts3t-s or $t$ for $j\in\{0,1,2d,2o,3x,4\}$.
Our induction step has the following pattern: For each $i\in\{0,1,2d,2o,3x,4\}$, we develop a set of properties $\mathscr{P}(i,j)$ such that each $\mathbb{F}\in{\Xi}_{i}(s,t)$ satisfies one of the properties. Each property $\mathscr{P}(i,j)$ identifies a minimal collection of sets included in $\mathbb{F}$ which we call an anchor.
For example, we have a set of six properties $\mathscr{P}(0,j)$ ($1\leq j\leq 5$) such that every set system $\mathbb{F}$ in ${\Xi}_{0}$ satisfies one of the six properties. $\mathscr{P}(0,1)$ states that there exists a pair of elements that appears in three distinct sets. If $\mathbb{F}\in{\Xi}_{0}$ satisfies this property, we know that $\mathbb{F}$ includes the anchor $\mathbb{A}=\{\{a,b,c\},\{a,b,d\},\{a,b,e\}\}$ for some distinct elements $a,b,c,d,$ and $e$.
Since every $t$-transversal of $\mathbb{F}$ is an extension of some transversal of the anchor, we enumerate the transversals of the anchor and for each such transversal we bound the number of extensions that lead to $t$-transversals. For enumerating the transversals of the anchor, We select certain elements (which we call core elements) that appear in the anchor and enumerate all transversals of the anchor based on the inclusion/exclusion of the core elements and any elements which are forced as a consequence. Let $\gamma$ be one such transversal and assume that it includes $j$ elements and excludes $e$ elements. The transversal $\gamma$ induces a restricted system $\mathbb{F}^{\prime}$ of sets. Specifically, $\mathbb{F}^{\prime}$ is obtained by eliminating all sets from $\mathbb{F}$ containing any included element of $\gamma$ and by deleting any excluded element from each set of $\mathbb{F}$. It is clear that $\mathbb{F}^{\prime}$ is a (tj)(t-j)-threshold 33-system over njen-j-e elements and its deficit is 3tn2j+e3t-n-2j+e. We also ensure that $\mathbb{F}^{\prime}\in{\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$. We now use induction to bound the number of (tj)(t-j)-transversals of $\mathbb{F}^{\prime}$.
There is another wrinkle to the induction step. While our discussion focused on a minimal anchor, the actual choice of the anchor depends on the occurrence pattern of core elements in sets outside of the anchor. We thus deal with several cases and in each case we work with an anchor which is obtained by extending a minimal anchor.
To continue with our example of the property $\mathscr{P}(0,1)$, we select $a$ and $b$ as our core variables. If there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set, we stick with the anchor $\mathbb{A}$. If not, we extend the anchor $\mathbb{A}$ to $\mathbb{A}^{\prime}$ include the additional sets in which $a$ or $b$ appear by themselves without the other element.
In each case, we enumerate all transversals of the anchor by including/excluding the core variables. The transversals (specified by the inclusion/exclusion of core elements), the resulting restricted set systems of $\mathbb{F}$ together with their parameters, and the inductive bounds of the number of transversals of the restricted systems are represented in a table as shown below for the anchor $\mathbb{A}$ when there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set. Without loss of generality (due to symmetry between $a$ and $b$), we assumed that $a$ does not appear in any set of $\mathbb{F}$ in which $b$ does not appear.
The columns with the label Core elements contain entries indicating whether the core elements are included in or excluded from the transversals of the anchor, $1$ signals inclusion and 0 signals exclusion. The column labeled Forced elements lists elements which are forced to be included or excluded as a consequence of the inclusion/exclusion of the core elements. The columns under the label Subsystem parameters describe the values of the parameters $n^{\prime}$, $s^{\prime}$, and $t^{\prime}$ as well as the type of the resulting restricted system. Each row (except the last row) represents a transversal of the anchorḞor each such row, the columns labeled Inductive bound contains the inductive upper bounds on the number of $t^{\prime}$-transversals for the restricted subsystem expressed as a fraction of the claimed upper bound on the number of $t$-transversals for $\mathbb{F}$ depending on the parity of $s$. In our example, we bound the number of transversals as a fraction of $(\frac{2}{3})^{\frac{s}{2}}3^{t}$ for even $s$. For odd $s$, we bound the number of transversals as a fraction of $(\frac{7}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$. The last row contains the sum of the fractions for even and odd parities of $s$. The induction step is successful if the fraction add up to less than or equal to $1$.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
0 1 n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
0 0 $c=d=e=1$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{12}$ $\frac{1}{12}$
Total $\frac{31}{36}$ $\frac{79}{84}$
Note 4.2.
To ensure that each restricted system belongs to ${\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$, we apply induction in the order the properties are stated.
In the next six sections, we present the proof of Theorem 4.1.
4.2 Proof of the upper bounds for $\Phi_{0}(s,t)$
We cover $\mathbb{F}\in{\Xi}_{0}(s,t)$ with the following properties.
$\mathscr{P}(0,1)$:
There exists a pair of elements that appears in at least three distinct $3$-sets.
$\mathscr{P}(0,2)$:
There exists an element that appears in at least three distinct $3$-sets.
$\mathscr{P}(0,3)$:
There exists a pair of elements that appears in two distinct $3$-sets.
$\mathscr{P}(0,4)$:
There exists a triangle configuration.
$\mathscr{P}(0,5)$:
There exists an $\mathbf{\exists}$} configuration.
$\mathscr{P}(0,6)$:
There exists an element that appears only in one set.
Lemma 4.3.
If a $3$-set system $\mathbb{F}$ does not satisfy the conditions $\mathscr{P}(0,2),\mathscr{P}(0,3)$ and $\mathscr{P}(0,4)$, then the following property holds: Let $\{a,b,c\}$ be any set in the system. For $x,y\in\{a,b,c\}$ and $x\neq y$, if there exists sets $S_{x}$, $S_{y}$ where $S_{x}$ contains $x$ and $S_{y}$ contains $y$ and $S_{x}$ and $S_{y}$ are distinct from $\{a,b,c\}$, then $S_{x}$ and $S_{y}$ are disjoint.
Lemma 4.4.
For every $\mathbb{F}\in{\Xi}_{0}$, one of the properties $\mathscr{P}(0,1)$ through $\mathscr{P}(0,6)$ holds.
Proof.
If a $3$-set system $\mathbb{F}$ does not satisfy the condition $\mathscr{P}(0,1)$, then every element appears at most twice. If, in addition, the property $\mathscr{P}(0,3)$ does not hold, then by lemma˜4.3, either there exists an $\mathbf{\exists}$} configuration or at least one element appear in exactly one set of $\mathbb{F}$. ∎
Note 4.5.
$\mathscr{P}(0,1)$ is redundant in the sense that every $\mathbb{F}\in{\Xi}_{0}$ satisfies one of the properties, $\mathscr{P}(0,i)$ for $2\leq i\leq 6$. However, we prioritize the application of induction based on $\mathscr{P}(0,1)$ to avoid dealing with restricted systems of three $2$-sets with a common element.
4.2.1 $\mathscr{P}(0,1)$ – There exists a pair of elements that appears in at least three distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,1)$. Let the distinct elements $a$ and $b$ appear together in the sets $\{a,b,c\},\{a,b,d\}$ and $\{a,b,e\}$ where $c$, $d$, and $e$ are distinct elements. Either one of $\{a,b\}$ is such that it always appears in a set together with the other element in $\{a,b\}$ or not. We consider each case in turn and work with the corresponding anchors.
4.2.1.1 $a$ always appears with $b$.
Without loss of generality (due to symmetry between $a$ and $b$), assume that $a$ always appears together with $b$.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
0 1 n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
0 0 $c=d=e=1$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{12}$ $\frac{1}{12}$
Total $\frac{31}{36}$ $\frac{79}{84}$
4.2.1.2 $a$ and $b$ each appear in a set without the presence of the other element.
We now extend the anchor by including sets containing either $a$ or $b$ without the other element being present. Observe that when $a=0$ and $b=1$ in the following table, we guaranteed the existence of at least one $2$-set in the resulting restricted system.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 0 $\frac{1}{2}$ $\frac{1}{2}$
0 1 n2n-2 t1t-1 s1s-1 11 $\frac{1}{3}$ $\frac{5}{14}$
0 0 $c=d=e=1$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{12}$ $\frac{1}{12}$
Total $\frac{11}{12}$ $\frac{79}{84}$
4.2.2 $\mathscr{P}(0,2)$ – There exists an element that appears in at least three distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,2)$. Let $a$ be an element that appears in three sets $\{a,b,c\}$, $\{a,d,e\}$ and $\{a,f,g\}$ of $\mathbb{F}$ where $\{b,c\}$, $\{d,e\}$ and $\{f,g\}$ are distinct and do not have a common element.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 0 $\frac{1}{2}$ $\frac{1}{2}$
0 n1n-1 $t$ $s+1$ 33 $\frac{1}{2}$ $\frac{1}{2}$
Total 11 11
4.2.3 $\mathscr{P}(0,3)$ – There exists a pair of elements that appears in two distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,3)$. Assume that $\mathbb{F}$ has a pair $\{a,b\}$ of elements which appears in exactly two sets $\{a,b,c\}$ and $\{a,b,d\}$ since otherwise if $\{a,b\}$ were to appear in three sets, we would have applied induction using $\mathscr{P}(0,1)$. We also know that $a$ (and $b$) cannot appear in any other set since otherwise we would have applied induction using $\mathscr{P}(0,2)$.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
0 n1n-1 $t$ $s+1$ $2o$ $\frac{5}{9}$ $\frac{4}{7}$
Total $\frac{17}{18}$ 11
4.2.4 $\mathscr{P}(0,4)$ – There exists a triangle configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,4)$. Let $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ be a triangle configuration in $\mathbb{F}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$. Moreover, we assume that $a$, $b$, and $c$ do not appear in any other set of $\mathbb{F}$.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 0 0 $e=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
0 11 0 $f=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
0 11 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
11 11 $c=e=f=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
0 11 11 $a=d=f=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
11 11 $b=e=d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
Total 11 11
4.2.5 $\mathscr{P}(0,5)$ – There exists an $\mathbf{\exists}$} configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,5)$. Let $\{a,b,c\}$ be the spine of the $\mathbf{\exists}$} configuration, and $\{a,d,e\},\{b,f,g\}$ and $\{c,h,i\}$ are its arms where the set $\{d,e,f,g,h,i\}$ is disjoint from $\{a,b,c\}$. We assume $a,b$, and $c$ do not appear in any other sets.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 0 0 n3n-3 t1t-1 $s$ $2d$ $\frac{25}{108}$ $\frac{5}{21}$
0 11 0 n3n-3 t1t-1 $s$ $2d$ $\frac{25}{108}$ $\frac{5}{21}$
0 0 11 n3n-3 t1t-1 $s$ $2d$ $\frac{25}{108}$ $\frac{5}{21}$
0 11 11 $f=g=h=i=0$ n7n-7 t2t-2 $s+1$ 11 $\frac{2}{27}$ $\frac{5}{63}$
11 0 11 $d=e=h=i=0$ n7n-7 t2t-2 $s+1$ 11 $\frac{2}{27}$ $\frac{5}{63}$
11 11 0 $d=e=f=g=0$ n7n-7 t2t-2 $s+1$ 11 $\frac{2}{27}$ $\frac{5}{63}$
11 11 11 $d=e=f=g=h=i=0$ n9n-9 t3t-3 $s$ 0 $\frac{1}{27}$ $\frac{1}{27}$
Total $\frac{103}{108}$ $\frac{187}{189}$
4.2.6 $\mathscr{P}(0,6)$ – There exists an element that appears only in one set
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,6)$. Let $a$ be such that $a$ appears exactly once in $\mathbb{F}$ in the set $\{a,b,c\}$. We consider two cases: Each of $b$ and $c$ appear in a different set or one of them does not appear in any other set.
4.2.6.1 Each of $b$ and $c$ appear in a different set:
Assume that $\mathbb{F}$ contains $\{a,b,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $d,e,f$ and $g$ are distinct and $a$ does not appear anywhere else.
Core elements Forced elements Subsystem parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 0 0 n3n-3 t1t-1 $s$ $2d$ $\frac{25}{108}$ $\frac{5}{21}$
0 11 0 n3n-3 t1t-1 $s$ 11 $\frac{5}{18}$ $\frac{2}{7}$
0 0 11 n3n-3 t1t-1 $s$ 11 $\frac{5}{18}$ $\frac{2}{7}$
0 11 11 d=e=f=g=0 n7n-7 t2t-2 $s+1$ 0 $\frac{7}{81}$ $\frac{2}{21}$
Total $\frac{283}{324}$ $\frac{19}{21}$
4.2.6.2 One of $\{b,c\}$ does not appear in any other set:
Assume that $\mathbb{F}$ contains $\{a,b,c\}$ and $a$ and $b$ do not appear anywhere else. Either $c$ does not appear in any other set or there is a set of the form $\{c,d,e\}$.
Core elements Forced elements Subsystem parameters Inductive bound
$c$ $a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 $a=b=0$ n3n-3 t1t-1 $s$ 0 $\frac{1}{3}$ $\frac{1}{3}$
0 11 0 n3n-3 t1t-1 $s$ 0 $\frac{1}{3}$ $\frac{1}{3}$
0 0 11 n3n-3 t1t-1 $s$ 0 $\frac{1}{3}$ $\frac{1}{3}$
Total 11 11
When $c=0$, the number of transversals when $c$ appears in another set is dominated by the number of transversals when $c$ does not appear in any other set.
4.3 Proof of the Upper bounds for $\Phi_{1}(s,t)$
Let $\Xi$} the family of $3$-set systems that contain a $2$-set. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n$ elements such that $\mathbb{F}=\{a,b\}\cup\mathbb{F}^{\prime}$ with threshold $t$ and deficit $s$. The following properties determine the subclasses.
$\mathscr{P}(1,1)$:
At least one variable in $\{a,b\}$ appears exactly once.
$\mathscr{P}(1,2)$:
At least one of the elements in $\{a,b\}$ appears at least twice.
4.3.1 At least one variable in $\{a,b\}$ appears exactly once  – $\mathscr{P}(1,1)$
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
0 11 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
Total $\frac{14}{15}$ 11
4.3.2 At least one of the elements in $\{a,b\}$ appears at least twice  – $(\neg\mathscr{P}(1,1))\land\mathscr{P}(1,2)$
In this case, $\mathbb{F}$ has a set of the form $\{a,c,d\}$ where $\{c,d\}$ is disjoint from $\{a,b\}$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 0 $\frac{3}{5}$ $\frac{7}{12}$
0 11 n2n-2 t1t-1 s1s-1 11 $\frac{2}{5}$ $\frac{5}{12}$
Total 11 11
4.4 Proof of the Upper bounds for $\Phi_{2o}(s,t)$
We now prove the upper bounds for $3$-set systems which contain two overlapping $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with two overlapping $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the overlapping $2$-sets $\{a,b\}$ and $\{a,c\}$.
We first deal with the simpler case of odd $s$. We will show that the number of $t$-transversals is at most $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Odd $s$
11 n1n-1 t1t-1 s2s-2 $\frac{7}{10}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{3}{10}$
Total 11
For the rest of the section, let $s$ be even. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{2}{3})(\frac{2}{3})^{\frac{s}{2}}3^{t}$. For this purpose, we consider the following subclasses of $\Xi$}.
$\mathscr{P}(2o,1)$:
$a$ appears three times.
$\mathscr{P}(2o,2)$:
Either $b$ or $c$ appears four times.
$\mathscr{P}(2o,3)$:
One of $b$ or $c$ appears only once.
$\mathscr{P}(2o,4)$:
There are two sets containing both $b$ and $c$.
$\mathscr{P}(2o,5)$:
There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
$\mathscr{P}(2o,6)$:
There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
$\mathscr{P}(2o,7)$:
For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
$\mathscr{P}(2o,8)$:
One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
$\mathscr{P}(2o,9)$:
There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$.
$\mathscr{P}(2o,10)$:
There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together.
Lemma 4.6.
If $\mathbb{F}$ satisfies $\neg(\bigwedge_{j=1}^{3}\mathscr{P}(2o,j))$, then $a$ appears exactly twice and $b$ and $c$ appear at least two times and at most three times in $\mathbb{F}$. Moreover, the following scenarios cover all the situations if one of $b$ or $c$ were to appear three times.
  1. 1.
    There are two sets containing both $b$ and $c$.
  2. 2.
    There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
  3. 3.
    There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
  4. 4.
    For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
  5. 5.
    One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
Lemma 4.7.
If $\mathbb{F}$ satisfies $\neg\bigwedge_{j=1}^{8}\mathscr{P}(2o,j)$, then $a$, $b$ and $c$ appear exactly twice.
4.4.1 $a$ appears three times  – $\mathscr{P}(2o,1)$
Without of loss of generality, assume that $a$ appears in $\{a,d,e\}$ in addition to appearing in $\{a,b\}$ and $\{a,c\}$ where neither $d$ nor $e$ are in $\{b,c\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 n1n-1 t1t-1 s2s-2 $\frac{3}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 1 $\frac{1}{4}$
Total 11
4.4.2 Either $b$ or $c$ appears four times  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2o,j))\land\mathscr{P}(20,2)$
Without of loss of generality, assume that $b$ appears in $\{a,b\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{b,h,i\}$ where none of $\{d,e\}$, $\{f,g\}$ and $\{h,i\}$ is equal to $\{a,c\}$.
We consider two cases:
4.4.2.1 Case 1:
There is no element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 n1n-1 t1t-1 s2s-2 11 $\frac{5}{8}$
0 $a=1$ n2n-2 t1t-1 s1s-1 33 $\frac{3}{8}$
Total 11
4.4.2.2 Case 2:
There is an element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$. Let $d=f=h$
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 n1n-1 t1t-1 s2s-2 11 $\frac{5}{8}$
0 0 $a=d=f=h=1$ n6n-6 t4t-4 s6s-6 0 $\frac{1}{16}$
0 11 $a=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{47}{48}$
4.4.3 One of $b$ or $c$ appears only once  – $(\neg\bigvee_{j=1}^{2}\mathscr{P}(2o,j))\land\mathscr{P}(2o,3)$
Assume that $b$ appears only once.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 $\frac{7}{12}$
0 11 $c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
Total $\frac{7}{8}$
4.4.4 There are two sets containing both $b$ and $c$  – $(\neg\bigvee_{j=1}^{3}\mathscr{P}(2o,j))\land\mathscr{P}(2o,4)$
Without of loss of generality, assume we have two sets $\{b,c,d\}$ and $\{b,c,e\}$ where $d\neq e$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 $d=e=1$ n5n-5 t3t-3 s4s-4 $\frac{1}{8}$
11 0 11 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
11 11 0 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
Total 11
4.4.5 There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$  – $(\neg\bigvee_{j=1}^{4}\mathscr{P}(2o,j))\land\mathscr{P}(2o,5)$
Without of loss of generality, assume we have a set $\{b,c,d\}$. Furthermore, assume that there is another set $\{b,e,f\}$ where neither $e$ nor $f$ is equal to $c$ and a set $\{c,g,h\}$ where neither $g$ nor $h$ is equal to $b$. We consider two cases:
4.4.5.1 $d\not\in\{e,f,g,h\}$:
$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 $a=1$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 11 $a=1$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 0 $a=d=1$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
11 11 $a=0$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.5.2 $d\in\{e,f,g,h\}$:
Assume that $d=e$. In this case, we have the sets $\{a,b\},\{a,c\},\{b,c,d\}$, $\{b,d,f\}$ and $\{c,g,h\}$.
$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 $a=1,d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
11 11 $a=0$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
0 $a=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
Total $\frac{11}{12}$
4.4.6 There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$   – $(\neg\bigvee_{j=1}^{5}\mathscr{P}(2o,j))\land\mathscr{P}(2o,6)$
Assume that there are sets $\{b,d,e\}$, $\{c,d,e\}$, and $\{b,f,g\}$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{23}{24}$
4.4.7 For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$  – $(\neg\bigvee_{j=1}^{6}\mathscr{P}(2o,j))\land\mathscr{P}(2o,7)$
Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$, $\{c,d^{\prime},e^{\prime}\}$ and $\{c,f^{\prime},g^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$ are distinct.
We consider three cases:
4.4.7.1 Case 1:
There is no element common to three or more of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 n3n-3 t1t-1 $s$ 33 $\frac{7}{24}$
11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.7.2 Case 2:
There is an element common to exactly three of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$. Assume without loss of generality, $d=f=d^{\prime}$ where $d$ is distinct from $f^{\prime}$ and $g^{\prime}$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
11 0 0 11 n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.7.3 Case 3:
There is an element common to exactly four of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$. Assume without loss of generality, $d=f=d^{\prime}=f^{\prime}$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=g^{\prime}=1$ n8n-8 t5t-5 s7s-7 0 $\frac{1}{24}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.8 One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element  – $(\neg\bigvee_{j=1}^{7}\mathscr{P}(2o,j))\land\mathscr{P}(2o,8)$
Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{c,d^{\prime},e^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$ are distinct.
We consider two cases:
Case 1: There is no element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 n3n-3 t1t-1 $s$ 33 $\frac{7}{24}$
11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 $2o$ $\frac{5}{36}$
11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{35}{36}$
Case 2: There is an element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$. Assume $d=f=d^{\prime}$
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 $2o$ $\frac{5}{36}$
11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{35}{36}$
4.4.9 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$  – $(\neg\bigvee_{j=1}^{8}\mathscr{P}(2o,j))\land\mathscr{P}(2o,9)$
Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\}$, and $\{b,c,d\}$ where $d$ is neither $b$ nor $c$. We consider two cases.
4.4.9.1 $d$ appears uniquely:
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
11 0 11 $d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
11 11 $c=d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
0 $b=c=1,d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
Total 11
4.4.9.2 $d$ does not appear uniquely:
There is another clause $\{d,f,g\}$ where $f$ and $g$ are distinct from $b$ and $c$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
11 0 11 $d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
11 11 $c=d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.10 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together  – $(\neg\bigvee_{j=1}^{9}\mathscr{P}(2o,j))\land\mathscr{P}(2o,10)$
Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $\{d,e,f,g\}\cap\{b,c\}=\varnothing$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 11 $\frac{1}{2}$
11 11 $c=d=e=0$ n5n-5 t2t-2 s1s-1 11 $\frac{1}{6}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{23}{24}$
4.5 Proof of the Upper bounds for $\Phi_{2d}(s,t)$
We now prove the upper bounds for $3$-set systems which contain two disjoint $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with two disjoint $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 4$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains two disjoint $2$-sets $\{a,b\}$ and $\{c,d\}$. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{5}{6})^{2}(\frac{2}{3})^{\frac{s}{2}}3^{t}$ when $s$ is even and by $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$ when $s$ is odd. We cover $\Xi$} based on the following properties.
$\mathscr{P}(2d,1)$:
One of $\{a,b,c,d\}$ appears 3 times.
$\mathscr{P}(2d,2)$:
There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$.
$\mathscr{P}(2d,3)$:
A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$.
$\mathscr{P}(2d,4)$:
One of $\{a,b,c,d\}$ appears exactly once.
$\mathscr{P}(2d,5)$:
There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
Lemma 4.8.
If $\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j)$, then $a$ and $b$ must each appear in another set which contains neither $c$ not $d$ and the two sets cannot have more than one common element.
Proof.
Under the hypothesis, it must be
  1. 1.
    none of $\{a,b,c,d\}$ appears more than twice,
  2. 2.
    none of $\{a,b,c,d\}$ appears exactly once, and
  3. 3.
    no pair from $\{a,b\}\times\{c,d\}$ appears in any set.
Since $a$ and $b$ must participate in two sets, let there be two sets $\{a,e,f\}$ and $\{b,g,h\}$ which are distinct from $\{a,b\}$ and $\{c,d\}$. It must be $\{e,f,g,h\}\cap\{a,b,c,d\}=\varnothing$ since no pair from $\{a,b\}\times\{c,d\}$ appears in any set. Moreover, $\{e,f\}\neq\{g,h\}$ since no pair occurs with one of $\{a,b,c,d\}$ more than once. ∎
4.5.1 One of $\{a,b,c,d\}$ appears 3 times  – $\mathscr{P}(2d,1)$
Assume that $\mathbb{F}$ contains an element in $\{a,b,c,d\}$, say $a$, which appears in three sets. Let $\{a,e,f\}$ and $\{a,g,h\}$ be the other two sets in which $a$ appears. Neither $\{e,f\}$ nor $\{g,h\}$ is equal to $\{c,d\}$.
We consider two cases.
4.5.1.1 Case 1:
There is no common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
0 11 n2n-2 t1t-1 s1s-1 33 $\frac{9}{25}$ $\frac{7}{20}$
Total $\frac{24}{25}$ $\frac{19}{20}$
4.5.1.2 Case 2:
There is a common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$. Let $e=f=d$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
0 11 0 $f=h=c=1$ n6n-6 t4t-4 s6s-6 0 $\frac{3}{50}$ $\frac{7}{120}$
0 11 11 n3n-3 t2t-2 s3s-3 0 $\frac{7}{25}$ $\frac{3}{10}$
Total $\frac{47}{50}$ $\frac{23}{24}$
4.5.2 There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2d,j))\land\mathscr{P}(2d,2)$
   Assume without loss of generality that $\mathbb{F}$ contains a set of the form $\{a,c,e\}$ where $e$ is distinct from $b$ and $d$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
0 1 n2n-2 t1t-1 s1s-1 22 overlapping $\frac{2}{5}$ $\frac{2}{5}$
Total 11 11
4.5.3 A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$  – $(\neg\bigwedge_{j=1}^{2}\mathscr{P}(2d,j))\land\mathscr{P}(2d,3)$
Assume that there are two sets $\{e,f,g\}$ and $\{e,f,h\}$ containing the pair $\{e,f\}$ where $g,h\in\{a,b,c,d\}$. Since each of $\{a,b,c,d\}$ appear at most twice, we know that neither $e$ nor $f$ are in $\{a,b,c,d\}$. There are two cases.
4.5.3.1 $g=a$ and $h=c$:
$a$ $c$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 $d=1,b=0$ n4n-4 t2t-2 s2s-2 1 $\frac{1}{5}$ $\frac{1}{5}$
0 11 $b=1,d=0$ n4n-4 t2t-2 s2s-2 1 $\frac{1}{5}$ $\frac{1}{5}$
0 0 $b=d=1$ n4n-4 t2t-2 s2s-2 1 $\frac{1}{5}$ $\frac{1}{5}$
11 11 n2n-2 t2t-2 s4s-4 0 $\frac{9}{25}$ $\frac{7}{20}$
Total $\frac{24}{25}$ $\frac{19}{20}$
4.5.3.2 $g=a$ and $h=b$:
$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 0 11 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
11 0 1 0 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
0 11 0 11 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
0 11 11 0 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
11 11 0 0 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
Total 11 11
4.5.4 One of $\{a,b,c,d\}$ appears exactly once  – $(\neg\bigwedge_{j=1}^{3}\mathscr{P}(2d,j))\land\mathscr{P}(2d,4)$
Assume that $a$ is unique.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 11 $\frac{12}{25}$ $\frac{1}{2}$
0 1 n2n-2 t1t-1 s1s-1 11 $\frac{12}{25}$ $\frac{1}{2}$
Total $\frac{24}{25}$ 11
4.5.5 There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$  – $(\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j))\land\mathscr{P}(2d,5)$
We have that every one of $\{a,b,c,d\}$ appears exactly twice in $\mathbb{F}$. Assume we have sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
11 1 $e=f=g=h=0$ n5n-5 t2t-2 s1s-1 11 $\frac{4}{25}$ $\frac{1}{6}$
Total $\frac{24}{25}$ 11
4.6 Proof of the Upper bounds for $\Phi_{3}(s,t)$
We now prove the upper bounds for $3$-set systems which contain three $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with three $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the $2$-sets $\{a,b\}$, $\{c,d\}$, and $\{e,f\}$ where $a,b,c,d,e$ and $f$ are not necessarily distinct. We regard the $2$-sets as edges of a graph, consider several cases based on the structure of the graph and prove the bounds in each case.
4.6.1 The three 2-sets form a path of length 3
Consider the case where $F$ contains three $2$-sets that form a path of length $3$. Let $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$ be the $2$-sets that form a path (of length $3$). We cover the family of such set systems by the following sub families.
$\mathscr{P}(3,1)$:
At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
$\mathscr{P}(3,2)$:
Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$
Fact 4.9.
Either $\mathscr{P}(3,1)$ or $\mathscr{P}(3,2)$ holds.
4.6.1.1 At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.  – $\mathscr{P}(3,1)$
Assume without loss of generality that $b$ appears in no other set.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
Total 11 11
4.6.1.2 Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$  – $\mathscr{P}(3,2)$
Assume without loss of generality that $b$ appears in another set which is distinct from the sets in the path. Let $\{b,e,f\}$ such a set. Note the $e$ and $f$ are distinct from $a$ and $c$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
1 $\{c,d\}$ remains n1n-1 t1t-1 s2s-2 11 $\frac{5}{7}$ $\frac{2}{3}$
1 0 1 $\{e,f\}$ remains n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
Total 11 $\frac{17}{18}$
4.6.2 The three 2-sets form a triangle
Let $\mathbb{F}$ be a set system with three 2-sets that form a triangle (cycle of length 3) configuration. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
At least one of $\{a,b,c\}$ is is no other set outside the triangle.
$\mathscr{P}(3,2)$:
At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set.
$\mathscr{P}(3,3)$:
Each of $\{a,b,c\}$ are in at least two other sets outside the triangle.
Lemma 4.10.
One of $\mathscr{P}(3,1),\mathscr{P}(3,2),\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
Proof.
In this case, by $\neg\mathscr{P}(3,1)$ $a,b,c$ are each in at least one clause besides the triangle. By $\neg\mathscr{P}(3,2)$, then, we know none of them is in exactly one clause besides the triangle, as all are in at least one. This means each is in at least two additional clauses, implying $\mathscr{P}(3,3)$. By contradiction, this therefore cannot occur. ∎
4.6.2.1 At least one of $\{a,b,c\}$ is is no other set outside the triangle  – $\mathscr{P}(3,1)$
We must select at least two elements of the triangle, otherwise either one of the clauses would be unsatisfied. Assume without loss of generality that $a$ is in no other set outside of the triangle. In this case, not all the elements in the triangle can be set to 1 simultaneously.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 1 0 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
Total 11 11
4.6.2.2 At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set  – $\mathscr{P}(3,2)$
Assume without loss of generality that $a$ is in exactly in one set $\{a,d,e\}$ outside the triangle where $d$ and $e$ are distinct from $b$ and $c$. Note that this set must not contain $b$ or $c$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
1 0 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
1 1 0 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
1 1 1 $d=e=0$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{7}$ $\frac{7}{54}$
Total 11 $\frac{26}{27}$
4.6.2.3 Each of $\{a,b,c\}$ are in at least two other sets outside the triangle  – $\mathscr{P}(3,3)$
The two clauses attached to any triangle member must not be redundant with each other, and they cannot contain other triangle variables without being redundant.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
1 0 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
1 1 0 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
1 1 1 n3n-3 t3t-3 s6s-6 0 $\frac{3}{14}$ $\frac{7}{36}$
Total $\frac{13}{14}$ $\frac{8}{9}$
4.6.3 There exists a 2-set that is disjoint from the other two 2-sets
Let $\mathbb{F}$ be a set system with three 2-sets that where one of the 2-sets is disjoint from the remaining 2-set. Without loss of generality, let this isolated 2-set be $\{a,b\}$. Let $\{c,d\}$ and $\{e,f\}$ be the other two sets.
We will first deal with the cases when $s$ is odd. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
$s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap.
$\mathscr{P}(3,2)$:
$s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$.
$\mathscr{P}(3,3)$:
$s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$.
Lemma 4.11.
One of $\mathscr{P}(3,1),\mathscr{P}(3,2)$ and $\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
4.6.3.1 $s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap  – $\mathscr{P}(3,1)$
If $\{c,d\}$ and $\{e,f\}$ overlap, the following calculation applies.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{5}{9}$
0 1 n2n-2 t1t-1 s1s-1 $2o$ $\frac{4}{9}$
Total 11
4.6.3.2 $s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,2)$
We take advantage of the fact that $a$ appears uniquely since $a$ and $b$ cannot be simultaneously $1$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
Total $\frac{25}{27}$
4.6.3.3 $s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
In this case, no single added 2-clause can possibly make the three form a star, as the existing two don’t share any points of overlap. We can therefore use the standard type 3 bound.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2d$ $\frac{5}{9}$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
Total $\frac{17}{18}$
We now deal with the case that $s$ is even. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
$s$ is even and either $a$ or $b$ appear in no other set.
$\mathscr{P}(3,2)$:
$s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$.
$\mathscr{P}(3,3)$:
$s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
$\mathscr{P}(3,4)$:
$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
$\mathscr{P}(3,5)$:
$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
$\mathscr{P}(3,6)$:
$s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists.
Lemma 4.12.
For $i\in[6]$, one of $\mathscr{P}(3,i)$ always holds for every set system in $\Xi$}.
Proof.
Let $F\in{\Xi}$ be arbitrary and assume none of $\mathscr{P}(3,i)$ holds for any choice of isolated 2-set in the three 2-set system.
Since $\mathscr{P}(3,1)$ and $\mathscr{P}(3,2)$ don’t hold, $a$ and $b$ appear in at least one other clause, and all such clauses don’t contain any of $c,d,e,f$. Notice that amongst $c,d,e,f$, by non-redundancy, either there exist $3$ distinct variables or $4$. If there were $3$, then we see that $\mathscr{P}(3,3)$ must have been satisfied. Hence, it must be the case that $c,d,e,f$ are all distinct.
This means that the three 3-sets in $F$ must be disjoint. Therefore by symmetry, instead of letting $\{a,b\}$ be the isolated 2-set, we can relabel variables and let any of $\{c,d\}$ or $\{e,f\}$ be the isolated 2-set and check for these properties. Hence, we can conclude that none of $a,b,c,d,e,f$ appears in any other set with each other (this is ruled out already for say $a,b$ by non-redundancy and remaining by negation of $\mathscr{P}(3,2)$ and symmetry). Also, we must have that each of $a,b,c,d,e,f$ must appear in at least one more 3-set. Since $\mathscr{P}(3,5)$ does not hold, it must be that both $a,b$ appear in exactly one other clause and by symmetry that each of $c,d,e,f$ appear in exactly one other clause. Also, since $\mathscr{P}(3,4)$ does not hold, we conclude that if the unique other 3-set where $a$ appears is say $\{a,x,y\}$ then it must be the case that there exists 3-set $\{c,x,y\}$. Symmetrically applying this argument to all pairs, we infer that case $\mathscr{P}(3,6)$ must hold for $F$, which is a contradiction. ∎
4.6.3.4 $s$ is even and either $a$ or $b$ appear in no other set  – $\mathscr{P}(3,1)$
Without loss of generality let this be $a$. We take advantage of the fact that $a$ appears uniquely.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
1 0 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
0 1 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
Total $\frac{20}{21}$
4.6.3.5 $s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$  – $\mathscr{P}(3,2)$
Without loss of generality let this be $a$ and $c$. Let this set be $\{a,c,x\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
0 1 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$
0 1 0 $d=x=1$ n5n-5 t3t-3 s4s-4 11 $\frac{5}{42}$
Total 11
4.6.3.6 $s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
Without loss of generality let $v_{ab}=a$ and assume $c=f$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{4}{7}$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
Total 11
4.6.3.7 $s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,4)$
Without loss of generality let $v_{ab}=a$ and $v_{cdef}=c$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
0 1 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$
0 1 0 1 n4n-4 t2t-2 s2s-2 33 $\frac{1}{6}$
Total 11
4.6.3.8 $s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,5)$
Without loss of generality let $v_{ab}=a$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
0 1 n2n-2 t1t-1 s1s-1 44 $\frac{17}{42}$
Total 11
4.6.3.9 $s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists  – $\mathscr{P}(3,6)$
We directly compute:
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
1 0 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
1 1 $x=y=0,c=d=e=f=1$ n8n-8 t6t-6 s10s-10 0 $\frac{1}{56}$
Total $\frac{7}{8}$
4.7 Proof of the Upper Bounds for $\Phi_{4}(s,t)$
Let $F=F_{1}F_{2}$ for some monotone $3$-CNF $F_{2}$, where $F_{1}$ is composed of four distinct 2-clauses and one of them is $(ab)$. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties. (Note we only care about odd values of $s$, as even values are handled by the three 2-clause case.)
$\mathscr{P}(4,1)$:
No two two-clauses intersect.
$\mathscr{P}(4,2)$:
Some variable is in exactly two 2-clauses.
$\mathscr{P}(4,3)$:
Some variable is in at least three 2-clauses.
Lemma 4.13.
Some $\mathscr{P}(4,j)$ must hold in a formula with at least have four two-clauses.
Proof.
Proceed by contradiction, assuming $\neg\bigwedge_{j=1}^{3}\mathscr{P}(4,j)$. In this case, by $\neg\mathscr{P}(4,1)$ some two-clause intersects another. This means some variable is in more than one 2-clause. By $\neg\mathscr{P}(4,2)$, it must be in at least three 2-clauses, implying $\mathscr{P}(4,3)$. ∎
4.7.0.1 No two two-clauses intersect  – $\mathscr{P}(4,1)$
In this case, no three of the 2-clauses form a star, as they don’t overlap at all. We can therefore use the standard type 3 bound.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 33 $\frac{9}{17}$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{7}{17}$
Total $\frac{16}{17}$
4.7.0.2 Some variable is in exactly two 2-clauses  – $\mathscr{P}(4,2)$
Without loss of generality, assume $a$ is in both $ab$ and $ac$. Note $b$ and $c$ can be in as many 2-clauses as desired.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 22 $\frac{10}{17}$
0 1 $c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{6}{17}$
Total $\frac{16}{17}$
4.7.0.3 Some variable is in at least three 2-clauses  – $\mathscr{P}(4,3)$
Without loss of generality, assume $a$ is in $ab$, acac, and $ad$. It may be in as many other 2-clauses as desired, as may $b$, $c$, and $d$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 0 $\frac{14}{17}$
0 1 $c=d=1$ n4n-4 t3t-3 s5s-5 0 $\frac{3}{17}$
Total 11
Appendix A Optimality of Building Blocks
A.0.1 ${\mathbb{T}}_{5}^{3}$ is $2$-extremal
Theorem A.1.
${\mathbb{T}}_{5}^{3}$ is $2$-extremal 33-system over $5$ elements.
Proof.
Let $\mathbb{H}$ be $3$-system on $5$ elements with transversal number at least $2$. We show that $\theta(\mathbb{H})\leq 7$. We assume that $\tau(\mathbb{H})=2$ since otherwise $\theta_{2}(\mathbb{H})=0$.
Case 1: Assume that $\mathbb{H}$ contains a singleton set $\{x\}$. Every transversal of $\mathbb{H}$ must contain $x$ and thus $\theta_{2}(\mathbb{H})=\theta_{1}(\mathbb{H}^{x})$. Since $\mathbb{H}^{x}$ is a 3-system on $4$ vertices and $\tau(\mathbb{H}^{x})\geq 1$, we have $\theta_{2}(\mathbb{H})=\theta_{1}(\mathbb{H}^{x})\leq\Theta(4,1,3)=3$.
Case 2: Assume that $\mathbb{H}$ contains a $2$-set $\{x,y\}$. $\theta_{2}(\mathbb{H})=\theta_{1}(\mathbb{H}^{x})+\theta_{2}(\mathbb{H}^{\neg x})$. $\theta_{1}(\mathbb{H}^{x})\leq\Theta(4,1,3)=3$ since $\mathbb{H}^{x}$ has $4$ elements. If $x$ is not in a transversal, then the transversal must contain $y$. $\mathbb{H}^{\neg x,y}$ is a $3$-system on $3$ vertices with transversal number at least 1. We thus have $\theta_{2}(\mathbb{H})\leq\Theta(4,4,3)+\Theta(3,1,3)=3+3<7$.
Case 3: Assume that $\mathbb{H}$ is a $3$-uniform set system and $\tau(\mathbb{H})=2$. $\mathbb{H}$ must have at least three $3$-sets in its collection. We consider two cases:
Case 3.1: $\mathbb{H}$ contains two sets of the form $\{x,y,u\}$ and $\{x,y,v\}$ where $u$ and $v$ are distinct. Consider the following mutually exclusive and exhaustive assignments to $x$ and $y$: 1. $x=1$, 2. $x=0,y=1$ and 3. $x=0,y=0$. In the first case, $\theta_{1}(\mathbb{H}^{x})\leq\Theta(4,1.3)=3$. In the second case, $\theta_{1}(\mathbb{H}^{\neg x,y})\leq\Theta(3,1.3)=3$. In the third case, we also have $u=v=1$ and hence the number of $2$-transversals is at most 1. Combining the conclusions of the cases, we get that $\theta_{2}(\mathbb{H})\leq 7$.
Case 3.2: $\mathbb{H}$ contains two sets of the form $\{x,y,z\}$ and $\{x,u,v\}$ where $u,v,x$ and $y$ are distinct. Consider the two cases: 1. $x=1$, and 2. $x=0$. In the first case, $\theta_{1}(\mathbb{H}^{x})\leq\Theta(4,1.3)=3$. In the second case, $\mathbb{H}^{\neg x}$ contains two disjoint $2$-sets which lets us conclude that $\theta_{2}(\mathbb{H}^{\neg x})\leq 4$. Combining the conclusions of the cases, we get that $\theta_{2}(\mathbb{H})\leq 7$. ∎
A.0.2 ${\mathbb{T}}_{6}^{3}$ is $3$-extremal
Theorem A.2.
${\mathbb{T}}_{6}^{3}$ is a $3$-extremal 33-system on $6$ vertices.
Proof.
Let $\mathbb{H}$ be a $3$-graph on $6$ vertices with $\tau(\mathbb{H})\geq 3$. We will show that $\theta(\mathbb{H})_{3}\leq 14$. We assume that $\tau(\mathbb{H})=3$ since otherwise $\theta(\mathbb{H})_{3}=0$.
Case 1: Assume that $\mathbb{H}$ contains a singleton set $\{x\}$. Every transversal of $\mathbb{H}$ must contain $x$ and thus $\theta_{3}(\mathbb{H})=\theta_{2}(\mathbb{H}^{x})$. Since $\mathbb{H}^{x}$ is a 3-graph on $5$ vertices and $\tau(\mathbb{H}^{x})\geq 2$, we have $\theta_{3}(\mathbb{H})\leq\Theta(5,2,3)=7$.
Case 2: Assume that $\mathbb{H}$ contains a $2$-set $\{x,y\}$. $\theta_{3}(\mathbb{H})=\theta_{2}(\mathbb{H}^{x})+\theta_{3}(\mathbb{H}^{\neg x})$. $\theta_{2}(\mathbb{H}^{x})\leq\Theta(5,2,3)$ since $\mathbb{H}^{x}$ has $5$ elements. If $x$ is not in a transversal, then the transversal must contain $y$. $\mathbb{H}^{\neg x,y}$ is a graph on $4$ vertices with transversal number at least 2. We thus have $\theta_{3}(\mathbb{H})\leq\Theta(5,2,3)+\Theta(4,2,3)=7+6<14$.
Case 3: Assume that $\mathbb{H}$ is a $3$-uniform set system and $\tau(\mathbb{H})=3$. It must be that $m(\mathbb{H})\geq 6$ which implies that there exists a an element $x$ which appears in at least $3$ sets. We have $\theta_{3}(\mathbb{H})=\theta_{2}(\mathbb{H}^{x})+\theta_{3}(\mathbb{H}^{\neg x})$. Since $\mathbb{H}^{x}$ has $5$ vertices, $\theta_{2}(\mathbb{H}^{x})\leq\Theta(5,2,3)=7$.
We now show that $\theta_{3}(\mathbb{H}^{\neg x})\leq 7$. $\mathbb{H}^{\neg x}$ has at least three distinct $2$-sets and $\tau(\mathbb{H}^{\neg x})\geq 3$. Select any three distinct $2$-sets from among these. At least two of these sets must have an element $y$ in common since $\mathbb{H}^{\neg x}$ has at most $5$ elements. We consider two sub cases.
Case 3.1: $y$ is common to all three 2-sets. Let $\{y,u\},\{y,v\}$ and $\{y,w\}$ be the three setss where $u,v$ and $w$ are distinct. $\theta_{3}(\mathbb{H}^{\neg x})=\theta_{2}(\mathbb{H}^{\neg x,y})+\theta_{3}(\mathbb{H}^{\neg x,\neg y})$. $\theta_{2}(\mathbb{H}^{\neg x,y})$ is at most $6$ since $\mathbb{H}^{\neg x,y}$ is a $3$-system on $4$ vertices with transversal number at least 2. $\theta_{3}(\mathbb{H}^{\neg x,\neg y})=1$ since every $3$-transversal of $\mathbb{H}^{\neg x,\neg y}$ must include $u,v,$ and $w$. This shows that $\theta_{3}(\mathbb{H}^{\neg x})\leq 7$ and hence $\theta_{3}(\mathbb{H})\leq 14$ in this case.
Case 3.2: Assume that $y$ is common to exactly two of the three $2$-sets. Let $\{y,u\}$ and $\{y.v\}$ be these two sets where $u\neq v$. Let $\{a,b\}$ be the remaining $2$-set which may contain $u$ or $v$. We will show $\theta(\mathbb{H}^{\neg x,y})_{2}\leq 5$ and $\theta_{3}(\mathbb{H}^{\neg x,\neg y})\leq 2$ which together imply that $\theta_{3}(\mathbb{H}^{\neg x})\leq 7$.
$\mathbb{H}^{\neg x,y}$ has at most $4$ elements and transversal number at least 2. If $\mathbb{H}^{\neg x,y}$ has no $3$-sets, then it must have at least one more $2$-set in addition to $\{a,b\}$ which implies that it has at most 44 22-transversals.
If $\mathbb{H}^{\neg x,y}$ has a $3$-set, assume that $a$ is in a $3$-set (without loss of generality). We have $\theta_{2}(\mathbb{H}^{\neg x,y})=\theta_{1}(\mathbb{H}^{\neg x,y,a})+\theta_{2}(\mathbb{H}^{\neg x,y,\neg a})$. The number of $1$-transversals of $\mathbb{H}^{\neg x,y,a}$ is at most three since there are at most 3 elements in $\mathbb{H}^{\neg x,y,a}$. $\theta_{2}(\mathbb{H}^{\neg x,y,\neg a})\leq 2$ since every $2$-transversal of $\mathbb{H}^{\neg x,y,\neg a}$ must include $b$ and there are at most two choices for the remaining element. In total we have at most 55 22-transversals in $\mathbb{H}^{\neg x,y}$ in this case.
The number of $3$-transversals of $\mathbb{H}^{\neg x,\neg y}$ is at most two. These $3$-transversals must contain $u$ and $v$ and one of the two remaining vertices.
In conclusion, we have $\theta_{3}(\mathbb{H})\leq 14$ in each case. ∎

### v2

@@
 1 Introduction
+What is the largest number of minimum weight satisfying assignments of an $n$-variate monotone 3-CNF with no satisfying assignment of Hamming weight less than $n/2$?
+We show a tight upper bound of $6^{n/4}$.
+A construction matching this bound is given by dividing the variables into disjoint blocks of size 4, and for each block including all the monotone 3-clauses.
+Our proof in fact gives a deterministic algorithm that enumerates these solutions.
+This problem is closely related to the complexity of Majority function and $k$-SAT.
+Circuit complexity.
 Ajtai [1] and Furst, Saxe, and Sipser [2] showed that the Parity function, which is easily computable in linear time, cannot be computed by constant-depth circuits of polynomial size.
@@
 $\Sigma^{k}_{3}$ lower bounds.
-Paturi, Pudlák, and Zane [13] (henceforth called PPZ) showed that Parity requires $\Sigma^{k}_{3}$ circuits of size $\Omega(2^{n/k})$.
+Paturi, Pudlák, and Zane [13] (hereafter called PPZ) showed that Parity requires $\Sigma^{k}_{3}$ circuits of size $\Omega(2^{n/k})$.
 In fact, this results in an unrestricted depth-3 lower bound of $\Omega(n^{1/4}2^{\sqrt{n}})$.
 Both of these lower bounds are known to be tight.
-Paturi, Pudlák, Saks, and Zane [12] (henceforth called PPSZ) extended the techniques of [13] and showed that recognizing the codewords of a good error-correcting code such as BCH requires $\Sigma^{k}_{3}$ circuits of size $2^{cn/k}$ for a universal constant $c>1$.
+Paturi, Pudlák, Saks, and Zane [12] (hereafter called PPSZ) extended the techniques of [13] and showed that recognizing the codewords of a good error-correcting code such as BCH requires $\Sigma^{k}_{3}$ circuits of size $2^{cn/k}$ for a universal constant $c>1$.
 This remains the best known lower bound for computing any explicit function by $\Sigma^{k}_{3}$ circuits.
@@
 Non-trivial $k$-SAT algorithms.
-An algorithm solving $k$-SAT is non-trivial if it has running time $2^{(1-\epsilon_{k})n}$ for some $\epsilon>0$ which we call savings.
+An algorithm solving $k$-SAT is non-trivial if it has running time $2^{(1-\epsilon_{k})n}$ for some $\epsilon_{k}>0$ which we call savings.
 PPZ and PPSZ used their combinatorial arguments to give non-trivial $k$-SAT algorithms.
 PPSZ currently holds the record as the best known $k$-SAT algorithm with savings $c/k$ for a universal constant $c>1$ [14].
-The Super Strong Exponential Time Hypothesis (SSETH) states that such savings, that is $\Theta(1/k)$, are best possible [18].
-We know that indeed this is the case for PPSZ [14].
+The Super Strong Exponential Time Hypothesis (SSETH) states that $\Theta(1/k)$ savings are best possible [17].
 Depth-3 complexity of Majority function.
@@
 They pointed out that determining $\Sigma^{k}_{3}$ complexity of Majority is non-trivial even for $k=3$, and proposed this as a first step towards settling the depth-3 complexity of Majority.
-Results of Lecomte, Ramakrishnan, and Tan [11] can be interpreted as proving the conjecture under the assumption that each depth-2 subcircuits depends on at most $k$ input bits.
+Results of Lecomte, Ramakrishnan, and Tan [11] can be interpreted as proving the conjecture under the assumption that each depth-2 sub-circuits depends on at most $k$ input bits.
 Similar to PPZ and PPSZ, determining the depth-3 complexity is closely tied to $k$-SAT.
 Gurumukhani et al.
-[6] showed that proving the conjectured lower bound for $\Sigma^{k}_{3}$ via a particular approach which they called local enumeration, would break SSETH.
+[6] showed that proving the conjectured lower bound for $\Sigma^{k}_{3}$ via a particular approach, which they called local enumeration, would break SSETH.
 Local enumeration.
 Enum($k$, $t$) is defined as follows.
-For a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF $F$ with no satisfying assignment of Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$.
+For a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF $\mathbb{F}$ with no satisfying assignment of Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$.
 Analogously, we can define a restriction of this problem to monotone formulas which we denote by Enum+($k$, $t$).
@@
 1.1 Our contribution
-Our main result is an optimal lower bound on the size of any monotone $\Sigma^{3}_{3}$ circuit computing Majority matching the upper bounds conjectured in the literature (see e.g. [4]).
-Theorem 1.2.
-$\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq 2^{O(n\log(k)/k)}$.
-In particular, $\operatorname{Size}^{3}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq(2/6^{1/4})^{n}\simeq 1.277^{n}$.
-Furthermore, there are monotone circuits that achieve this bound.
-We optimally solve Enum+($3$, $n/2$) to derive our lower bound
+Our main result is an optimal lower bound on the size of any monotone $\Sigma^{3}_{3}$ circuit computing Majority matching the upper bounds conjectured in the literature (see e.g.
+[4]) which state $\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq 2^{O(n\log(k)/k)}$ and, in particular, $\operatorname{Size}^{3}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq(2/6^{1/4})^{n}\simeq 1.277^{n}$.
+We prove our lower bound by optimally solving Enum+($3$, $n/2$).
 Theorem 1.2 (Main result).
@@
 Consequently, $\operatorname{Size}^{+3}_{3}(\operatorname{\mathsf{Maj}}_{n})\geq(2/6^{1/4})^{n-o(n)}\simeq 1.277^{n}$.
-While we do not claim that our techniques are new, we believe that our approach to prove matching lower bounds for depth-3 circuits holds promise.
-The first element of our approach is the top-down technique: if a $\Sigma\Pi\Sigma_{3}^{k}$ circuit computes a Boolean function $f:\{0,1\}^{n}\mapsto\{0,1\}$ and $s$ is the number of its depth-2 circuits, one of the depth-2 circuits which is a $k$-CNF must accept at least $|f^{-1}(1)|/s$ members of $f^{-1}(1)$.
+The optimal algorithm for Enum+($3$, $t$) for $t\leq n/2$ is in turn obtained by showing via a constructive argument that extremal monotone $3$-CNF consistent with Majority function (explained below) cannot accept more than $6^{n/4}$ transversals (satisfying assignments) of Hamming weight $\lfloor\frac{n}{2}\rfloor$.
+While we do not claim that our techniques are fundamentally new, we believe that our approach to prove matching lower bounds for depth-3 circuits holds promise.
+The first element of our approach is the top-down technique: if a $\Sigma_{3}^{k}$ circuit computes a Boolean function $f:\{0,1\}^{n}\mapsto\{0,1\}$ and $s$ is the number of its depth-2 circuits (equivalently fan-in of the top OR gate), one of the depth-2 circuits (which is a $k$-CNF) must accept at least $|f^{-1}(1)|/s$ members of $f^{-1}(1)$.
 This brings our focus to the study of extremal $k$-CNFs with respect to $f$.
-A $k$-CNF $F$ is extremal with respect to $f$ if ${\textsc{Sat}}(F)\subseteq f^{-1}(1)$ and $|{\textsc{Sat}}(F)|$ is the maximum among such $k$-CNF.
-Such extremal $k$-CNFs can also used for constructing depth-3 $\Sigma\Pi\Sigma_{3}^{k}$ circuits which are (almost) optimal depending on the symmetries of $f$.
-Thus, the study of $k$-CNF which are extremal with respect to a Boolean function is of fundamental concern.
-What do we know of $k$-CNFs which are extremal with respect to the majority function or more generally with respect to Boolean threshold functions?
-Our intuition (based on extensive provably optimal constructions for small $n$) suggested that a modularity principle might hold.
-We say that modularity principle holds for a Boolean function if an extremal $k$-CNF $F$ can be expressed as the conjunction of $k$-CNFs $F^{\prime}_{i}$ where $F^{\prime}_{i}$ is supported on a bounded number of variables and the supports of $F^{\prime}_{i}$ are disjoint.
-Although we lack a formal method to determine if the majority function is inherently modular, we used the modularity principle to guide a trial-and-error search for constructing putative extremal 3-CNFs of Boolean threshold functions (for thresholds less or equal to $n/2$).
-Our final step is to apply induction to prove the optimality of our constructions.
-Several ingredients helped the induction to succeed: 1.
-Correct guess (or construction) of optimal constructions for Boolean threshold function with threshold less or equal to $n/2$, 2.
-Construction of (optimal or near optimal) extremal $3$-CNFs which are further restricted in certain ways (for example, $3$-CNFs which are required to include a $2$-clause), and 3.
-identification of appropriate anchors to carry out the induction taking advantage of forced variables (this ingredient in explain in detail in Section 4.
-In a nutshell, starting with precisely optimal constructions played an essential role in the proving optimality of our constructions.
-2 Set Systems
-A set system $\mathbb{F}=(V,F)$ is a collection $F$ of sets over a set $V$ of elements.
-If $V$ is a set of vertices, then $\mathbb{F}$ is interpreted as a graph.
-If $V$ is as a set of Boolean variables, then $\mathbb{F}$ is interpreted as a monotone CNF.
-While we regard $\mathbb{F}$ as a set system most of the time, we invite the reader to apply other interpretations to gain additional viewpoints.
-When we refer to sets in $\mathbb{F}$, we mean the sets in $F$.
+A $k$-CNF $\mathbb{F}$ is extremal with respect to $f$ if ${\textsc{Sat}}(\mathbb{F})\subseteq f^{-1}(1)$ ($\mathbb{F}$ is consistent with $f$) and $|{\textsc{Sat}}(\mathbb{F})|$ is the maximum among such $k$-CNF.
+Such extremal $k$-CNFs can also used for constructing $\Sigma_{3}^{k}$ circuits (almost) optimally depending on the symmetries of $f$.
+Thus, the study of $k$-CNFs which are extremal with respect to a Boolean function is of fundamental concern.
+What do we know of $k$-CNFs which are extremal with respect to Majority function or more generally with respect to Boolean threshold functions?
+Our intuition (based on optimal constructions for small $n$) suggested that a modularity principle might hold for extremal $k$-CNF with respect to Majority.
+We say that the modularity principle holds for a Boolean function if one of its extremal $k$-CNF $\mathbb{F}$ can be expressed as the conjunction of $k$-CNFs $\mathbb{F}^{\prime}_{i}$ where $\mathbb{F}^{\prime}_{i}$ is supported on a bounded number of variables and the supports of $\mathbb{F}^{\prime}_{i}$ are disjoint.
+Although we lacked a formal method to determine if the Majority function is modular a priori, we used the modularity principle to guide our search for extremal 3-CNFs of Boolean threshold functions for small values of $n$ and then generalized the constructions for all $n$.
+Our final step is to apply induction to prove the optimality of our constructions. Several ingredients helped the induction to succeed:
+  1. 1.
+    Inductive structure in the constructions of extremal monotone $3$-CNFs for general and restricted $3$-CNFs (for example, $3$-CNFs which are required to include a $2$-clause),
+  2. 2.
+    Correct guesses of extremal (restricted and unrestricted) $3$-CNFs for Boolean threshold functions with threshold less or equal to $n/2$, and
+  3. 3.
+    Intricate division of formulas (which are consistent with threshold functions) into subclasses.
+In a nutshell, we used an approach of upper bounds guiding the lower bounds where guessing the optimal constructions played an essential role in the proving their optimality.
+Our approach to the proof leads to an algorithm to enumerate the transversals.
+2 Monotone CNFs and Transversals
+A CNF (for Conjunctive Normal Form) formula $\mathbb{F}=(V,F)$ is a collection $F$ of clauses (where each clause is a set of literals) over a set $V$ of Boolean variables.
+We assume without loss of generality that the literals in the clause correspond to distinct variables.
 Notation 2.1.
-We use $n(\mathbb{F})$ to denote the number of elements and $m(\mathbb{F})$ to denote the number of sets of $\mathbb{F}$.
-We use $k$-set to refer to a set with cardinality $k$.
-We extend the notation to $k$-clauses and $k$-edges.
-We often work with set systems that where the size of the sets is bounded.
-We say that a set system $\mathbb{F}=(V,F)$ is a $k$-system if each $C\in F$ has cardinality at most $k$.
-If every set in the set system is a $k$-set, then we call it a $k$-uniform set system.
-We now introduce the essential concept of a transversal of a set system.
+We use $n(\mathbb{F})$ to denote the number of variables and $m(\mathbb{F})$ to denote the number of clauses of $\mathbb{F}$.
+We use $k$-clause to refer to a clause with cardinality $k$.
+We extend the notation to $k$-sets and $k$-edges.
+We often work with CNFs that where the size of the clauses is bounded.
+We say that a CNF $\mathbb{F}=(V,F)$ is a $k$-CNF if each $C\in F$ has cardinality at most $k$.
+If every clause in the CNF is a $k$-clause, then we call it a $k$-uniform CNF.
+In this paper, we will only deal with monotone CNFs, where each clause consists of only positive literals.
+A clause which only contains positive literals is called a monotone clause.
+If $V$ is a set of vertices, then a monotone CNF $\mathbb{F}$ is interpreted as a hypergraph.
+If $V$ is a set of elements, then $\mathbb{F}$ is interpreted as a set system.
+While we speak in the language of CNFs $\mathbb{F}$, we invite the reader to apply other interpretations to gain additional viewpoints.
+When we refer to clauses in $\mathbb{F}$, we mean the clauses in $F$.
+We now introduce the concept of a transversal (or satisfying assignment) of a monotone CNF.
 Definition 2.2 (Transversals).
-Let $\mathbb{F}=(V,F)$ be a set system.
+Let $\mathbb{F}=(V,F)$ be a monotone CNF.
 An $\ell$-transversal of $\mathbb{F}$ is a subset of $V$ of size $\ell$ that intersects every member of $F$.
 We use transversal to refer to an $\ell$-transversal for some $\ell\geq 0$.
-Fact 2.3 (Critical sets of a transversal [13]).
-If $\gamma$ is a minimum-size transversal of a set system $\mathbb{F}$, for every $x\in\gamma$ there exists a set $S_{x}$ in $\mathbb{F}$ such that $S_{x}\cap\gamma=\{x\}$.
-We call $S_{x}$ a critical set for $x$ with respect to $\gamma$.
-We parameterize set systems by the minimum size of its transversals to arrive at the notion of $t$-threshold systems.
-Definition 2.4 ($t$-threshold set systems).
-A set system is called $t$-threshold if all its transversals have size at least $t$.
+Fact 2.3 (Critical clauses of a transversal [13]).
+If $\gamma$ is a minimum-size transversal of a monotone CNF $\mathbb{F}$, for every $x\in\gamma$ there exists a clause $C_{x}$ in $\mathbb{F}$ such that $C_{x}\cap\gamma=\{x\}$.
+We call $C_{x}$ a critical clause for $x$ with respect to $\gamma$.
+We parameterize monotone CNFs by the minimum size of its transversals to arrive at the notion of $t$-threshold systems.
+Definition 2.4 ($t$-threshold CNFs).
+A monotone CNF is called a $t$-threshold CNF if all its transversals have size at least $t$.
 We occasionally need the following concept.
 Definition 2.5 (Transversal number).
-The transversal number $\tau(\mathbb{H})$ of a set system $\mathbb{H}$ is the cardinality of its minimum-size transversals.
+The transversal number $\tau(\mathbb{F})$ of a monotone CNF $\mathbb{F}$ is the cardinality of its minimum-size transversals.
 Notation 2.6.
 We use $\gamma_{t}(\mathbb{F})$ to denote the set of $t$-transversals of $\mathbb{F}$ and $\theta_{t}(\mathbb{F})$ to denote the number of $t$-transversals of $\mathbb{F}$.
-Definition 2.7 (Extremal $t$-threshold $k$-set systems).
-We say that a $t$-threshold $k$-system $\mathbb{F}$ is extremal if $\theta_{t}(\mathbb{F})$ is the maximum over all $t$-threshold $k$-systems on $n(\mathbb{F})$ elements.
+Definition 2.7 (Extremal $t$-threshold $k$-CNFs).
+We say that a $t$-threshold monotone $k$-CNF $\mathbb{F}$ is extremal if $\theta_{t}(\mathbb{F})$ is the maximum over all $t$-threshold monotone $k$-CNFs over $n(\mathbb{F})$ variables.
 Definition 2.8.
-For $n,k\geq 1$ and $0\leq t\leq n$, we define $\Theta(n,t,k)$ to be equal to $\theta_{t}(\mathbb{H})$ where $\mathbb{H}$ is an extremal $t$-threshold $k$-system over $n$ elements.
-The following states that uniform set systems suffice as far as extremal $t$-threshold $k$-set systems are concerned.
+For $n,k\geq 1$ and $0\leq t\leq n$, we define $\Theta(n,t,k)$ to be equal to $\theta_{t}(\mathbb{H})$ where $\mathbb{H}$ is an extremal $t$-threshold monotone $k$-CNF over $n$ variables.
+Note that uniform $k$-CNFs suffice as far as extremal $t$-threshold monotone $k$-CNFs are concerned.
 Fact 2.9.
-For $t\leq n-k+1$ and for every $t$-threshold $k$-system, there is a $t$-threshold $k$-uniform set system which has the same set of $t$-transversals.
-For t=nk+it=n-k+i and $1\leq i\leq k$, the $t$-threshold (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is extremal.
+For $t\leq n-k+1$ and for every $t$-threshold monotone $k$-CNF, there is a $t$-threshold monotone $k$-uniform CNF which has the same set of $t$-transversals.
+For t=nk+it=n-k+i and $1\leq i\leq k$, the $t$-threshold monotone (ki+1)(k-i+1)-uniform CNF consisting of all (ki+1)(k-i+1)-sets is extremal.
 2.1 Circuits and Transversals
-$k$-systems play a key role in the (monotone) circuit complexity of Boolean functions.
-Our circuits are $\Sigma\Pi\Sigma_{k}$ circuits which are depth-3 circuits where the top gates is an $\mathbf{OR}$ gate and its depth-2 subcircuits are $k$-CNFs.
-The size of such a circuit is the number of its depth-2 subcircuits.
-For monotone $\Sigma\Pi\Sigma_{k}$ circuits, the depth-2 subcircuits are $k$-systems.
+Monotone $k$-CNFs play a key role in the monotone circuit complexity of Boolean functions.
+We consider $\Sigma_{3}^{k}$ circuits which are depth-3 circuits where the top gate is an $\mathbf{OR}$ gate and its depth-2 sub-circuits are $k$-CNFs.
+The size of such a circuit is the number of its depth-2 sub-circuits.
+For monotone $\Sigma_{3}^{k}$ circuits, the depth-2 sub-circuits are monotone $k$-CNFs.
 Definition 2.10.
-A Boolean function $f$ on $n$ variables $x_{1},\ldots,x_{n}$ is a $t$-threshold function for $0\leq t\leq n$ provided $f(x_{1},\ldots,x_{n})=1$ if and only if $\Sigma_{i}x_{i}\geq t$.
-If a $k$-system $\mathbb{F}$ is a depth-2 subcircuit of a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function, then $\mathbb{F}$ is a $t$-threshold system.
+On $n$ variables $x_{1},\ldots,x_{n}$, a $t$-threshold function $f$ for $0\leq t\leq n$ is a Boolean function with $f(x_{1},\ldots,x_{n})=1$ if and only if $\Sigma_{i}x_{i}\geq t$.
+If a monotone $k$-CNF $\mathbb{F}$ is a depth-2 sub-circuit of a monotone $\Sigma^{k}_{3}$ circuit that computes a $t$-threshold function, then $\mathbb{F}$ is a $t$-threshold CNF.
 Since no such $\mathbb{F}$ has more than $\Theta(n,t,k)$ $t$-transversals, we conclude the following.
 Fact 2.11.
-If a monotone $\Sigma\Pi\Sigma_{k}$ circuit computes a $t$-threshold function, then its size is at least $\binom{n}{t}/\Theta(n,t,k)$.
-Furthermore, it is known that we only need extremal $t$-threshold $k$-systems as depth-2 subcircuits to construct size-optimal (up to a polynomial factor) monotone $\Sigma\Pi\Sigma_{k}$ circuits for computing $t$-threshold functions.
-For every $t$-threshold $k$-system $\mathbb{F}$, it is known that there exists a $\Sigma\Pi\Sigma_{k}$ circuit of size ${\mathtt{poly}}(n)\binom{n}{t}/\theta(\mathbb{F})$ for computing a $t$-threshold function.
+If a monotone $\Sigma^{k}_{3}$ circuit computes a $t$-threshold function, then its size is at least $\binom{n}{t}/\Theta(n,t,k)$.
+Furthermore, it is known that we only need extremal $t$-threshold monotone $k$-CNFs as depth-2 sub-circuits to construct size-optimal (up to a polynomial factor) monotone $\Sigma^{k}_{3}$ circuits for computing $t$-threshold functions.
+For every $t$-threshold monotone $k$-CNF $\mathbb{F}$, it is known that there exists a $\Sigma^{k}_{3}$ circuit of size ${\mathtt{poly}}(n)\binom{n}{t}/\theta(\mathbb{F})$ for computing a $t$-threshold function.
 This can be shown by exploiting the symmetries of the threshold functions and monotonicity [4].
 Fact 2.12.
-There exists a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function whose size is at most ${\mathtt{poly}}(n)\binom{n}{t}/\Theta(n,t,k)$.
-2.2 Connection to Turán Systems
-The following definition is adapted from [16, 9].
-Definition 2.13 (Turán system and Turán number).
-For integers $n\geq\ell\geq k\geq 1$, a $k$-system over an $n$-element set is called an $(n,\ell,k)$-Turán system if every $\ell$-set contains a set of the system.
-The Turán number ${\Psi}(n,\ell,k)$ is defined as
-${\Psi}(n,\ell,k)=\max_{j\leq k}\min m(\mathbb{F}),$
-where the minimum is taken over all $(n,\ell,j)$-Turán systems $\mathbb{F}$.
-Fact 2.14.
-Assume that $2k\leq n-1$. A collection of $k$-sets of an $n$-element set is an $(n,k+1,k)$-Turán system if and only if it is a (nk)(n-k)-threshold $k$-set system.
-Proof.
-Let $\mathbb{F}$ be an $(n,k+1,k)$-Turán system.
-Since every $(k+1)$-set includes a $k$-set of $\mathbb{F}$, every (nk1)(n-k-1)-set fails to intersect some member of $\mathbb{F}$.
-Thus $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-uniform set system.
-Conversely, if $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-set system, then no (nk1)(n-k-1)-set is a transversal for $\mathbb{H}$.
-By considering complements, we conclude $\mathbb{F}$ must then be an $(n,k+1,k)$-Turán system.
-∎
-Furthermore, we have
-Fact 2.15.
-$\Theta(n,n-k,k)\;=\;\binom{n}{k}\;-\;{\Psi}(n,k+1,k).$
-2.3 Configurations
-To prove upper bounds on the number of $t$-transversals of a $t$-threshold system, our arguments will be based on whether certain configurations of sets are present in the set system.
-In the following, we catalog these configurations.
-Definition 2.16 (Triangle configuration).
-We say that a set system contains a triangle configuration if it contains three sets of the form $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
-Definition 2.17 ( $\mathbf{\exists}$} configuration).
-We say that a set system contains a $\mathbf{\exists}$} configuration if it contains four sets of the form $\{a,b,c\}$, $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ where the set $\{d,e,f,g,h,i\}$ is disjoint from the set $a,b,c$.
-We say that $\{a,b,c\}$ is the spine and $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ are the arms of the $\mathbf{\exists}$} configuration.
-We sometimes regard a collection of $2$-sets as a graph where each $2$-set is an edge and use graph-theoretic terminology to describe configurations.
-Definition 2.18 (Path configuration).
-We say that a collections of $2$-sets is a path of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{\ell}\}$ where $a_{0},a_{1},\ldots,a_{\ell}$ are distinct.
-Definition 2.19 (Cycle configuration).
-We say that a collection of $2$-sets is a cycle of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{0}\}$ where $a_{0},a_{1},\ldots,a_{\ell-1}$ are distinct.
-3 Constructions for Extremal Threshold System
-In this section, we construct $k$-systems over $n\geq 1$ elements for transversal number $0\leq t\leq\lfloor\frac{n}{2}\rfloor$.
+There exists a monotone $\Sigma^{k}_{3}$ circuit that computes a $t$-threshold function whose size is at most ${\mathtt{poly}}(n)\binom{n}{t}/\Theta(n,t,k)$.
+3 Constructions for Extremal Threshold CNFs
+In this section, we construct $k$-CNFs over $n\geq 1$ variables for transversal number $0\leq t\leq\lfloor\frac{n}{2}\rfloor$.
 In the next section, we argue that they are $t$-extremal.
@@
 Definition 3.1 (Clique).
-For $l\geq k\geq 1$, let ${\mathbb{K}}_{l}^{k}$ denote the $k$-system of all $k$-sets on $l$ elements.
-The following two cliques are widely used in our constructions: ${\mathbb{K}}_{3}^{3}=\{\{1,2,3\}\}$ is a $1$-threshold 33-system where $\theta_{1}({\mathbb{K}}_{3}^{3})=3$; ${\mathbb{K}}_{4}^{3}=\{\{1,2,3\},\{1,2,4\},\{1,3,4\},\{2,3,4\}\}$ is a $2$-threshold 33-system where $\theta_{2}({\mathbb{K}}_{4}^{3})=6$.
+For $l\geq k\geq 1$, let ${\mathbb{K}}_{l}^{k}$ denote the monotone $k$-CNF of all monotone $k$-clauses on $l$ variables.
+The following two cliques are widely used in our constructions: ${\mathbb{K}}_{3}^{3}=\{\{1,2,3\}\}$ is a $1$-threshold monotone $3$-CNF where $\theta_{1}({\mathbb{K}}_{3}^{3})=3$; ${\mathbb{K}}_{4}^{3}=\{\{1,2,3\},\{1,2,4\},\{1,3,4\},\{2,3,4\}\}$ is a $2$-threshold monotone $3$-CNF where $\theta_{2}({\mathbb{K}}_{4}^{3})=6$.
 Fact 3.2.
-${\mathbb{K}}_{l}^{k}$ is an extremal (lk+1)(l-k+1)-threshold $k$-uniform set system on $l$ elements where $n({\mathbb{K}}_{l}^{k})=l,$ $m({\mathbb{K}}_{l}^{k})=\binom{l}{k}$, $\tau({\mathbb{K}}_{l}^{k})=l-k+1,$ and $\Theta(l,l-k+1,k)=\theta_{l-k+1}({\mathbb{K}}_{l}^{k})=\binom{l}{l-k+1}$.
+${\mathbb{K}}_{l}^{k}$ is an extremal (lk+1)(l-k+1)-threshold $k$-uniform monotone CNF on $l$ variables where $n({\mathbb{K}}_{l}^{k})=l,$ $m({\mathbb{K}}_{l}^{k})=\binom{l}{k}$, $\tau({\mathbb{K}}_{l}^{k})=l-k+1,$ and $\Theta(l,l-k+1,k)=\theta_{l-k+1}({\mathbb{K}}_{l}^{k})=\binom{l}{l-k+1}$.
 3.2 Turán-style constructions
 Let $n\geq 3$.
-Constructions from here on are primarily limited to $k=3$.
-Consider the following (n,n3,3)(n,n-3,3)-Turán system ${\mathbb{T}}_{n}^{3}$ which is obtained by what we call a TuránTur\'{a}n-style construction [16].
-Partition $n$ into three parts where the parts have sizes as equal as possible and organize them in a circular fashion.
-${\mathbb{T}}_{n}^{3}$ consists of the following $3$-sets:
+We observe that $\Theta(n,t,3)$ is closely related to the hypergraph Turán problem.
+In particular, t=n3t=n-3 exactly captures the tetrahedron problem [15, 9].
+Consider the following (n3)(n-3)-threshold formula ${\mathbb{T}}_{n}^{3}$ which is obtained by what we call a TuránTur\'{a}n-style construction.
+Partition the $n$ variables into three parts where the parts have sizes as equal as possible and organize them in a circular fashion.
+${\mathbb{T}}_{n}^{3}$ consists of the following $3$-clauses:
   1. 1.
-    For each part, all 3-sets of the part
+    For each part, all 3-clauses of the part
   2. 2.
-    For each part, all 3-sets of the form where two elements come from the part and one element from the adjacent part in the circular order.
-Conjecture 3.3 (Turán).
-[…]
-The conjecture implies
-Fact 3.4.
-Conjecture 3.3 implies
-[…]
+    For each part, all 3-clauses of the form where two variables come from the part and one variable from the adjacent part in the circular order.
 We use the following Turán-style building blocks in our constructions.
 Example 3.3.
-${\mathbb{T}}_{5}^{3}=\{\{1,2,3\},\{3,4,5\},\{1,2,4\}\}$ is a $2$-threshold 33-uniform set system, $n({\mathbb{T}}_{5}^{3})=5$, $m({\mathbb{T}}_{5}^{3})=3$, $\tau({\mathbb{T}}_{5}^{3})=2$, and $\theta_{2}({\mathbb{T}}_{5}^{3})=7$.
+${\mathbb{T}}_{5}^{3}=\{\{1,2,3\},\{3,4,5\},\{1,2,4\}\}$ is a $2$-threshold 33-uniform monotone CNF, $n({\mathbb{T}}_{5}^{3})=5$, $m({\mathbb{T}}_{5}^{3})=3$, $\tau({\mathbb{T}}_{5}^{3})=2$, and $\theta_{2}({\mathbb{T}}_{5}^{3})=7$.
 Example 3.4.
 ${\mathbb{T}}_{6}^{3}=\{\{x_{0},y_{0},x_{1}\},\{x_{0},y_{0},y_{1}\},\{x_{1},y_{1},x_{2}\},\{x_{1},y_{1},y_{2}\},\{x_{2},y_{2},x_{0}\},\{x_{2},y_{2},y_{0}\}\}$ is a $3$-threshold 33-uniform set system, $n({\mathbb{T}}_{6}^{3})=6$, $m({\mathbb{T}}_{6}^{3})=6$, $\tau({\mathbb{T}}_{6}^{3})=3$, and $\theta_{3}({\mathbb{T}}_{6}^{3})=14$.
-Proofs of the following statement are provided in the appendix (Theorems A.1 and A.2).
-Fact 3.7.
+Proofs of the following statement are provided in the appendix (Theorems A.2 and A.3).
+Fact 3.5.
 $\Theta(5,2,3)$ $=\theta_{2}({\mathbb{T}}_{5}^{3})=7$
@@
 In this section, we provide constructions that work for all $n,k\geq 1$, $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$.
-We need the following notation to represent larger set systems composed of smaller set systems.
-If $\mathbb{F}_{1}=(V_{1},F_{1})$ and $\mathbb{F}_{2}=(V_{2},F_{2})$ are $k$-systems over disjoint sets of elements, the sum $\mathbb{F}_{1}+\mathbb{F}_{2}$ of $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ represents the $k$-system $\mathbb{F}=(V_{1}\cup V_{2},F_{1}\cup F_{2})$.
-For a $k$-system $\mathbb{F}$, $\ell\cdot\mathbb{F}$ represents the $k$-system obtained by summing $\ell$ disjoint copies of $\mathbb{F}$.
-Let $\mathbb{S}^{k}_{n,t}=t\cdot{\mathbb{K}}_{k}^{k}$. For $n\geq 1$ and $0\leq t\leq\lfloor n/k\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique optimal construction for any $k\geq 1$.
+We need the following notation to represent larger CNFs composed of smaller CNFs.
+If $\mathbb{F}_{1}=(V_{1},F_{1})$ and $\mathbb{F}_{2}=(V_{2},F_{2})$ are monotone $k$-CNFs over disjoint sets of variables, the sum $\mathbb{F}_{1}+\mathbb{F}_{2}$ of $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ represents the monotone $k$-CNF $\mathbb{F}=(V_{1}\cup V_{2},F_{1}\cup F_{2})$.
+For a monotone $k$-CNF $\mathbb{F}$, $\ell\cdot\mathbb{F}$ represents the monotone $k$-CNF obtained by summing $\ell$ disjoint copies of $\mathbb{F}$.
+Let $\mathbb{S}^{k}_{n,t}=t\cdot{\mathbb{K}}_{k}^{k}$. For $n\geq 1$ and $0\leq t\leq\lfloor n/k\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique extremal construction for any $k\geq 1$.
 Theorem 3.6 ([4]).
-For $k,n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{k}\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique (up to isomorphism) extremal $t$-threshold $k$-system on $n$ elements.
+For $k,n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{k}\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique (up to isomorphism) extremal $t$-threshold monotone $k$-CNF on $n$ variables.
 For t=nk+it=n-k+i where $1\leq i\leq k$, we have
-Fact 3.9.
-The (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is an extremal $t$-threshold $k$-system and $\Theta(n,n-k+i,k)=\binom{n}{n-k+i}$.
+Fact 3.7.
+The (ki+1)(k-i+1)-uniform monotone CNF consisting of all (ki+1)(k-i+1)-clauses is an extremal $t$-threshold monotone $k$-CNF and $\Theta(n,n-k+i,k)=\binom{n}{n-k+i}$.
 3.4 Constructions for $n\geq 1,k=3$ and $0\leq t\leq\lfloor\frac{n}{2}\rfloor$
@@
 They use ${\mathbb{K}}_{3}^{3}$, ${\mathbb{K}}_{4}^{3}$, ${\mathbb{T}}_{5}^{3}$, and ${\mathbb{T}}_{6}^{3}$ as building blocks.
-The following important reparameterization is useful to express our constructions.
+Our constructions vary based on the parity of 3tn3t-n, so the following reparameterization is useful.
 Definition 3.8.
-For $k=3$, we rewrite n=3tsn=3t-s where $s$ is called the deficit of the elements (variables or vertices) with respect to the threshold $t$.
-To emphasize the role of $s$, we specialize $\Theta(n,t,k)$ for $k=3$ and use the function $\Phi_{0}(s,t)$ to denote the maximum number of $t$-transversals of a $t$-threshold 33-system with deficit $s$.
+For $k=3$, we rewrite n=3tsn=3t-s where $s$ is called the deficit of the variables (elements or vertices) with respect to the threshold $t$.
+Our constructions and bounds depend on the parity of $s$.
+To emphasize the role of $s$, we specialize $\Theta(n,t,k)$ for $k=3$ and use the function $\Phi_{0}(s,t)$ to denote the maximum number of $t$-transversals of a $t$-threshold monotone $3$-CNF with deficit $s$.
 In other words, $\Phi_{0}(s,t)=\Theta(3t-s,t,3)$.
-Let $s\leq t$ and $t\geq 0$ (equivalently $n\geq 2t\geq 0$). Let $\mathbb{P}^{3}_{s,t}$ be the $3$-system on n=3tsn=3t-s elements defined as follows.
+Let $s\leq t$ and $t\geq 0$ (equivalently $n\geq 2t\geq 0$). Let $\mathbb{P}^{3}_{s,t}$ be the $3$-CNF over n=3tsn=3t-s variables defined as follows.
 […]
-Fact 3.11.
-$\mathbb{P}^{3}_{s,t}$ is a $t$-threshold 33-uniform set system.
-Fact 3.12.
+Fact 3.9.
+$\mathbb{P}^{3}_{s,t}$ is a $t$-threshold 33-uniform monotone CNF.
+Fact 3.10.
 […]
 Remark 3.11.
-${\mathbb{K}}_{3}^{3}+{\mathbb{T}}_{6}^{3}$ can be replaced ${\mathbb{K}}_{4}^{3}+{\mathbb{T}}_{5}^{3}$ with out affecting the number of minimum-size transversals.
-We summarize the results of this section by the following theorem which provides lower bounds for the maximum number of minimum-size transversals $\Phi(s,t)$ of $3$-systems over n=3tsn=3t-s elements with transversal number and deficit satisfying the conditions, $s\leq t$ and $t\geq 0$.
+${\mathbb{K}}_{3}^{3}+{\mathbb{T}}_{6}^{3}$ can be replaced with ${\mathbb{K}}_{4}^{3}+{\mathbb{T}}_{5}^{3}$ without affecting the number of minimum-size transversals.
+We summarize the results of this section by the following theorem.
+It provides lower bounds for the maximum number of minimum-size transversals $\Phi_{0}(s,t)$ of monotone $3$-CNFs over n=3tsn=3t-s elements with transversal number and deficit satisfying the conditions $s\leq t$ and $t\geq 0$.
 Theorem 3.12.
@@
 Proof.
-The first equation follows from [4]. The second equation is straightforward. The remaining two inequalities follow from Fact 3.12. ∎
+The first equation follows from [4]. The second equation is straightforward. The remaining two inequalities follow from Fact 3.10. ∎
 4 Upper Bounds on $\Phi_{0}(s,t)$
-In this section, we establish matching upper bounds on the number of transversals $\Phi_{0}(s,t)$ of extremal $t$-threshold 33-systems for $t\geq 0$ and $s\leq t$.
+In this section, we outline our plan to establish matching upper bounds on the number of transversals $\Phi_{0}(s,t)$ of extremal $t$-threshold monotone $3$-CNFs for $t\geq 0$ and $s\leq t$.
 We prove these bounds by induction on 3ts3t-s and $t$.
-As we do induction on $s$ and $t$, we have to deal with restricted set systems which motivates the following definitions.
-For $i\in\{1,2,3,4\}$, we say a $3$-system $\mathbb{F}$ is a type $i$ 33-system if it contains at least $i$ 22-sets.
-We further specialize the type 2 systems: $\mathbb{F}$ is a type $2o$ 33-system if it contains two $2$-sets which have exactly one element in common and a type $2d$ 33-system if it contains two $2$-sets which are disjoint.
-We also restrict type 3 systems as follows: $\mathbb{F}$ is a type $3x$ 33-system if it contains three $2$-sets which do not have an element in common.
-By extension, unrestricted $3$-systems are of type 0.
-For $i\in\{1,2d,2o,3x,4\}$, let
-  • ${\Xi}_{i}$ denote the family of $3$-system of type $i$,
-  • ${\Xi}_{i}(s,t)$ denote the family of $t$-threshold 33-system of type $i$ over n=3tsn=3t-s elements, and
-  • $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold 33-systems over 3ts3t-s elements.
+As we do induction on $n$ and $t$, we have to deal with restricted monotone CNFs, which motivates the following definitions.
+For $i\in\{1,2,3,4\}$, we say a monotone $3$-CNF $\mathbb{F}$ is a type $i$ monotone $3$-CNF if it contains at least $i$ monotone $2$-clauses.
+We further specialize the type 2 systems: $\mathbb{F}$ is a type $2o$ monotone $3$-CNF if it contains two monotone $2$-clauses which have exactly one variable in common and a type $2d$ 33-CNF if it contains two monotone $2$-clauses which are disjoint.
+As for type 3 CNFs, we only need to work with a restricted class of type 3 formulas, that is, formulas that do not contain three monotone $2$-clauses which have a variable in common.
+Since we do not consider other subtypes of type 3 CNFs, we will use type 3 CNFs to refer to the restricted class.
+By extension, unrestricted monotone $3$-CNFs are of type 0.
+For $i\in\{1,2d,2o,3,4\}$, let
+  • ${\Xi}_{i}$ denote the family of monotone $3$-CNFs of type $i$,
+  • ${\Xi}_{i}(s,t)$ denote the family of $t$-threshold monotone $3$-CNFs of type $i$ over n=3tsn=3t-s variables, and
+  • For $i\in\{0,1\}$, let $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold monotone $3$-CNFs over 3ts3t-s variables.
+For $i\in\{2d,2o,3,4\}$, let $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold monotone $3$-CNFs $F$ over 3ts3t-s variables where there is no pair of variables that appear in three clauses of $F$.
 We prove upper bounds on $\Phi_{0}(s,t)$ by simultaneously proving upper bounds on $\Phi_{i}(s,t)$ by induction on 3ts3t-s and $t$. Our theorem on upper bounds is presented below.
@@
 The following upper bounds hold for $\Phi_{i}(s,t)$ for $i\in\{1,2d,2o,3,4\}$.
-3-systems
+Monotone $3$-CNFs
 […]
-3-systems with at least one $2$-set
+Monotone $3$-CNFs with at least one $2$-clause
 […]
-3-systems with two $2$-sets which overlap in one element
+Monotone $3$-CNFs with two $2$-clauses which overlap in one variable
 […]
-3-systems with two $2$-sets which are disjoint
+Monotone $3$-CNFs with two $2$-clauses which are disjoint
 […]
-3-systems with three $2$-sets which do not have an element in common
+Monotone $3$-CNFs with three $2$-clauses which do not have an variable in common
 […]
-3-systems with four $2$-sets
+Monotone $3$-CNFs with four $2$-clauses
 […]
+Recall that we are primarily interested in the number of transversals of size $\lfloor n/2\rfloor$ in an $n$-vertex 3-graph with no transversal of smaller size.
+By setting $t=\lfloor n/2\rfloor$ and $s=t$, our bound follows.
+Corollary 4.2.
+$\Theta(n,\lfloor n/2\rfloor,3)\leq 6^{n/4}$.
 4.1 Outline of the proof
-The proof assumes that $t\geq 1$ as the theorem is trivially true otherwise.
-We divide the task of proving the upper bounds into six parts.
-In part $i\in\{0,1,2d,2o,3x,4\}$, we prove the upper bounds for $\Phi_{i}(s,t)$, the number of $t$-transversals for set systems in ${\Xi}_{i}(s,t)$ assuming the upper bounds for $\Phi_{j}(s,t)$ for smaller values of 3ts3t-s or $t$ for $j\in\{0,1,2d,2o,3x,4\}$.
-Our induction step has the following pattern: For each $i\in\{0,1,2d,2o,3x,4\}$, we develop a set of properties $\mathscr{P}(i,j)$ such that each $\mathbb{F}\in{\Xi}_{i}(s,t)$ satisfies one of the properties.
-Each property $\mathscr{P}(i,j)$ identifies a minimal collection of sets included in $\mathbb{F}$ which we call an anchor.
-For example, we have a set of six properties $\mathscr{P}(0,j)$ ($1\leq j\leq 5$) such that every set system $\mathbb{F}$ in ${\Xi}_{0}$ satisfies one of the six properties.
-$\mathscr{P}(0,1)$ states that there exists a pair of elements that appears in three distinct sets.
-If $\mathbb{F}\in{\Xi}_{0}$ satisfies this property, we know that $\mathbb{F}$ includes the anchor $\mathbb{A}=\{\{a,b,c\},\{a,b,d\},\{a,b,e\}\}$ for some distinct elements $a,b,c,d,$ and $e$.
-Since every $t$-transversal of $\mathbb{F}$ is an extension of some transversal of the anchor, we enumerate the transversals of the anchor and for each such transversal we bound the number of extensions that lead to $t$-transversals.
-For enumerating the transversals of the anchor, We select certain elements (which we call core elements) that appear in the anchor and enumerate all transversals of the anchor based on the inclusion/exclusion of the core elements and any elements which are forced as a consequence.
-Let $\gamma$ be one such transversal and assume that it includes $j$ elements and excludes $e$ elements.
-The transversal $\gamma$ induces a restricted system $\mathbb{F}^{\prime}$ of sets.
-Specifically, $\mathbb{F}^{\prime}$ is obtained by eliminating all sets from $\mathbb{F}$ containing any included element of $\gamma$ and by deleting any excluded element from each set of $\mathbb{F}$.
-It is clear that $\mathbb{F}^{\prime}$ is a (tj)(t-j)-threshold 33-system over njen-j-e elements and its deficit is 3tn2j+e3t-n-2j+e.
-We also ensure that $\mathbb{F}^{\prime}\in{\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$.
-We now use induction to bound the number of (tj)(t-j)-transversals of $\mathbb{F}^{\prime}$.
-There is another wrinkle to the induction step.
-While our discussion focused on a minimal anchor, the actual choice of the anchor depends on the occurrence pattern of core elements in sets outside of the anchor.
-We thus deal with several cases and in each case we work with an anchor which is obtained by extending a minimal anchor.
-To continue with our example of the property $\mathscr{P}(0,1)$, we select $a$ and $b$ as our core variables.
-If there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set, we stick with the anchor $\mathbb{A}$.
-If not, we extend the anchor $\mathbb{A}$ to $\mathbb{A}^{\prime}$ include the additional sets in which $a$ or $b$ appear by themselves without the other element.
-In each case, we enumerate all transversals of the anchor by including/excluding the core variables.
-The transversals (specified by the inclusion/exclusion of core elements), the resulting restricted set systems of $\mathbb{F}$ together with their parameters, and the inductive bounds of the number of transversals of the restricted systems are represented in a table as shown below for the anchor $\mathbb{A}$ when there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set.
-Without loss of generality (due to symmetry between $a$ and $b$), we assumed that $a$ does not appear in any set of $\mathbb{F}$ in which $b$ does not appear.
-The columns with the label Core elements contain entries indicating whether the core elements are included in or excluded from the transversals of the anchor, $1$ signals inclusion and 0 signals exclusion.
-The column labeled Forced elements lists elements which are forced to be included or excluded as a consequence of the inclusion/exclusion of the core elements.
-The columns under the label Subsystem parameters describe the values of the parameters $n^{\prime}$, $s^{\prime}$, and $t^{\prime}$ as well as the type of the resulting restricted system.
-Each row (except the last row) represents a transversal of the anchorḞor each such row, the columns labeled Inductive bound contains the inductive upper bounds on the number of $t^{\prime}$-transversals for the restricted subsystem expressed as a fraction of the claimed upper bound on the number of $t$-transversals for $\mathbb{F}$ depending on the parity of $s$.
-In our example, we bound the number of transversals as a fraction of $(\frac{2}{3})^{\frac{s}{2}}3^{t}$ for even $s$.
-For odd $s$, we bound the number of transversals as a fraction of $(\frac{7}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
-The last row contains the sum of the fractions for even and odd parities of $s$.
-The induction step is successful if the fraction add up to less than or equal to $1$.
-Core elements Forced elements Subsystem parameters Inductive bound
+The theorem trivially holds for $t=0$, so our proof focuses on the inductive step.
+For each $i$ and for all $j\in\{0,1,2d,2o,3,4\}$, we assume our upper bounds on $\Phi_{j}(s,t)$, the number of $t$-transversals of monotone CNFs in ${\Xi}_{j}(s,t)$, hold for all smaller values of 3ts3t-s or $t$.
+We use this to prove the upper bounds also hold for $\Phi_{i}(s,t)$.
+Our inductive step has the following pattern: For each $i\in\{0,1,2d,2o,3,4\}$, we develop a set of properties $\mathscr{P}(i,j)$ such that each $\mathbb{F}\in{\Xi}_{i}(s,t)$ satisfies at least one of the properties.
+Each property $\mathscr{P}(i,j)$ identifies a minimal collection of clauses included in $\mathbb{F}$, which we call an anchor.
+For example, we have a set of six properties $\mathscr{P}(0,j)$ ($1\leq j\leq 6$) such that every monotone CNF $\mathbb{F}$ in ${\Xi}_{0}$ satisfies at least one of the six properties.
+$\mathscr{P}(0,1)$ states that there exists a pair of variables that appears in three distinct clauses.
+If $\mathbb{F}\in{\Xi}_{0}$ satisfies this property, we know that $\mathbb{F}$ includes the clauses in $\mathbb{A}=\{\{a,b,c\},\{a,b,d\},\{a,b,e\}\}$ where the variables $a,b,c,d,$ and $e$ are distinct.
+We consider two cases to determine the anchor: either $a$ always appears together with $b$ in $\mathbb{F}$ or not.
+In the first case, we take $\mathbb{A}$ as our anchor.
+In the other case, our anchor will be $\mathbb{A}^{\prime}=\mathbb{A}\cup\{C\}$ where $C\in\mathbb{F}$ is a clause that contains $a$ but not $b$.
+Since every $t$-transversal of $\mathbb{F}$ is an extension of some transversal of the anchor, we enumerate all transversals of the anchor, and for each such transversal we bound the number of extensions that lead to $t$-transversals.
+To do so, we select certain variables (which we call core variables) that appear in the anchor and enumerate the transversals of the anchor by considering all settings (to 1/0) of the core variables depending on whether they are included in or excluded from a transversal and noting any variables which are forced as a consequence.
+Let $\gamma$ be one such transversal and assume that it includes $j$ variables and excludes $e$ variables in the anchor.
+The transversal $\gamma$ induces a restricted CNF $\mathbb{F}^{\prime}$.
+Specifically, $\mathbb{F}^{\prime}$ is obtained by eliminating all clauses from $\mathbb{F}$ containing any included variable of $\gamma$ and by deleting any excluded variable from each clause of $\mathbb{F}$.
+It is clear that $\mathbb{F}^{\prime}$ is a (tj)(t-j)-threshold monotone $3$-CNF over njen-j-e variables and its deficit is 3tn2j+e3t-n-2j+e.
+We also ensure that $\mathbb{F}^{\prime}\in{\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3,4\}$ so we can inductively bound the number of (tj)(t-j)-transversals by $\Phi_{i}(s+k-2j,t-j)$.
+We will continue with our example where $a$ always appears together with $b$.
+We use $\mathbb{A}$ as our anchor and $a$ and $b$ as core variables.
+When $a$ is set to $1$, we know $b$ must be set to 0 since if $b$ were set to $1$, $a$ would not have a critical clause.
+The resulting formula is of type 0 with $n^{\prime}=n-2$.
+Its threshold $t^{\prime}$ is at least t1t-1 and $s^{\prime}=s-1$.
+We now apply induction — if $s$ is even, $\Phi_{0}(s^{\prime},t^{\prime})=\Phi_{0}(s-1,t-1)=\frac{7}{18}\Phi_{0}(s,t)$, and if $s$ is odd, $\Phi_{0}(s^{\prime},t^{\prime})=\Phi_{0}(s-1,t-1)=\frac{3}{7}\Phi_{0}(s,t)$.
+When $a$ is set to 0, $b$ is set to either 0 or $1$.
+In either case, the resulting formula is of type 0.
+The first case results in exactly the same bounds as the case when $a$ is set to $1$.
+In the second case, we have to set $c$, $d$ and $e$ to $1$ so we get $n^{\prime}=n-5$, $t^{\prime}\geq t-3$ and s=s4s=s-4.
+We apply induction to get that the ratio of the number of transversals of the restricted formula to $\Phi_{0}(s,t)$ is at most $\frac{1}{12}$ regardless of the parity of $s$.
+The sum of the fractions is at most $1$, so our induction is successful.
+Our proof is presented in the form of a table as shown below.
+The columns with the label Core variables contain entries indicating whether the core variables are included in or excluded from the transversals of the anchor, $1$ signaling inclusion and 0 exclusion.
+The column labeled Forced variables lists variables which are forced to be included or excluded as a consequence of the setting of a subset of core variables.
+The columns under the label Restricted CNF parameters describe the values of the parameters and the type of the restricted CNF.
+Each row (except the last row) represents a transversal of the anchor.
+For each such row, the columns labeled Inductive bound contain the ratio of the inductive upper bounds to the claimed upper bound depending on the parity of $s$.
+The last row contains the sum of the fractions. The induction step is successful if the fractions add up to a total less than or equal to $1$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total $\frac{31}{36}$ $\frac{79}{84}$
-Note 4.2.
-To ensure that each restricted system belongs to ${\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$, we apply induction in the order the properties are stated.
-In the next six sections, we present the proof of Theorem 4.1.
+Note 4.3.
+To ensure that each restricted CNF belongs to ${\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3,4\}$, we apply induction in the order the properties are stated.
+Note 4.4.
+In the inductive step, we occasionally have situations in which the exact number of variables in the restricted formula is not guaranteed.
+In such cases, we inductively use $\Phi_{i}(s,t)\leq\Phi_{i}(s^{\prime},t)$ for every type $i$, $t$, and $s^{\prime}\leq s$.
+Note 4.5.
+In the inductive step, when the restricted formula has type 2 but it cannot be determined whether it is type $2o$ or $2d$, we observe the inductive hypothesis that $\Phi_{2o}(s,t)\leq\Phi_{2d}(s,t)$ and use the bound corresponding to type $2d$.
+Algorithmic aspect.
+It is easy to observe that our proof in fact gives a branching algorithm for enumerating the corresponding transversals.
+In each case, the proof identifies an anchor (a sub-formula), sets the core variables, and solves the restricted problems recursively.
+We will explicitly describe the algorithm in the full version of the paper.
+5 Conclusions
+Our results suggest several obvious open problems.
+We have only managed to prove optimal bounds for monotone $\Sigma_{3}^{3}$ for $t$-threshold functions for $t\leq n/2$.
+It would be desirable to extend these results for general $\Sigma_{3}^{3}$.
+It is perhaps more challenging to extend the current techniques for $\Sigma_{3}^{k}$ for $k\geq 4$ to settle the question regarding the complexity of Majority function.
+There are several obstacles.
+It is not clear whether the modularity principle holds for the Boolean threshold functions for $k$-CNF with $k\geq 4$.
+Even if it holds, it is not clear how to uniformly (with respect to $k$) find the building blocks.
+It would also be interesting to extend the construction for thresholds beyond $n/2$ even for monotone $\Sigma_{3}^{3}$.
+While we have certain guesses for the extremal constructions, proving their optimality may involve the resolution of famous hypergraph Turán problems.
+It is also interesting to find a shorter version (or a simpler alternative) of our proof.
+It is conceivable an intelligent computer search might be able to discover a shorter version of the current proof.
+Acknowledgments: We want to thank Pavel Pudlák for helpful discussions.
+6 Proof of theorem˜4.1
+In the rest of the paper, we present the proof of Theorem 4.1.
+6.1 Configurations
+To prove upper bounds on the number of $t$-transversals of a $t$-threshold monotone $k$-CNF, our arguments will be based on whether certain configurations of clauses are present in the CNF.
+In the following, we catalog these configurations for use in our proofs.
+Definition 6.1 (Triangle configuration).
+We say that a monotone CNF contains a triangle configuration if it contains three clauses of the form $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ where the clause $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
+Definition 6.2 ( $\mathbf{\exists}$} configuration).
+We say that a CNF contains a $\mathbf{\exists}$} configuration if it contains four clauses of the form $\{a,b,c\}$, $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ where the set $\{d,e,f,g,h,i\}$ is disjoint from the set $\{a,b,c\}$.
+We call $\{a,b,c\}$ the configuration’s spine and $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ its arms.
+We sometimes regard a collection of monotone $2$-clauses as a graph where each $2$-clause is an edge and use graph-theoretic terminology to describe configurations.
+Definition 6.3 (Path configuration).
+We say that a collections of monotone $2$-clauses is a path of length $\ell$ if it contains $\ell$ clauses of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{\ell}\}$ where $a_{0},a_{1},\ldots,a_{\ell}$ are distinct variables.
+Definition 6.4 (Cycle configuration).
+We say that a collection of $2$-clauses is a cycle of length $\ell$ if it contains $\ell$ clauses of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{0}\}$ where $a_{0},a_{1},\ldots,a_{\ell-1}$ are distinct variables.
 6.2 Proof of the upper bounds for $\Phi_{0}(s,t)$
@@
 $\mathscr{P}(0,1)$:
-There exists a pair of elements that appears in at least three distinct $3$-sets.
+There exists a pair of variables that appears in at least three distinct $3$-clauses.
 $\mathscr{P}(0,2)$:
-There exists an element that appears in at least three distinct $3$-sets.
+There exists a variable that appears in at least three distinct $3$-clauses.
 $\mathscr{P}(0,3)$:
-There exists a pair of elements that appears in two distinct $3$-sets.
+There exists a pair of variables that appears in two distinct $3$-clauses.
 $\mathscr{P}(0,4)$:
@@
 $\mathscr{P}(0,6)$:
-There exists an element that appears only in one set.
+There exists a variable that appears only in one clause.
 Lemma 6.5.
-If a $3$-set system $\mathbb{F}$ does not satisfy the conditions $\mathscr{P}(0,2),\mathscr{P}(0,3)$ and $\mathscr{P}(0,4)$, then the following property holds: Let $\{a,b,c\}$ be any set in the system.
-For $x,y\in\{a,b,c\}$ and $x\neq y$, if there exists sets $S_{x}$, $S_{y}$ where $S_{x}$ contains $x$ and $S_{y}$ contains $y$ and $S_{x}$ and $S_{y}$ are distinct from $\{a,b,c\}$, then $S_{x}$ and $S_{y}$ are disjoint.
+If a monotone $3$-CNF $\mathbb{F}$ does not satisfy the conditions $\mathscr{P}(0,2),\mathscr{P}(0,3)$ and $\mathscr{P}(0,4)$, then the following property holds: Let $\{a,b,c\}$ be any clause in the system.
+For $x,y\in\{a,b,c\}$ and $x\neq y$, if there exists clausess $S_{x}$, $S_{y}$ where $S_{x}$ contains $x$ and $S_{y}$ contains $y$ and $S_{x}$ and $S_{y}$ are distinct from $\{a,b,c\}$, then $S_{x}$ and $S_{y}$ are disjoint.
 Lemma 6.6.
@@
 Proof.
-If a $3$-set system $\mathbb{F}$ does not satisfy the condition $\mathscr{P}(0,1)$, then every element appears at most twice.
-If, in addition, the property $\mathscr{P}(0,3)$ does not hold, then by lemma˜4.3, either there exists an $\mathbf{\exists}$} configuration or at least one element appear in exactly one set of $\mathbb{F}$.
+If a monotone $3$-CNF $\mathbb{F}$ does not satisfy the condition $\mathscr{P}(0,2)$, then every variable appears at most twice.
+If, in addition, the property $\mathscr{P}(0,3)$ does not hold, then by lemma˜6.5, either there exists an $\mathbf{\exists}$} configuration or at least one variable appear in exactly one clause of $\mathbb{F}$.
 ∎
-Note 4.5.
+Note 6.7.
 $\mathscr{P}(0,1)$ is redundant in the sense that every $\mathbb{F}\in{\Xi}_{0}$ satisfies one of the properties, $\mathscr{P}(0,i)$ for $2\leq i\leq 6$.
-However, we prioritize the application of induction based on $\mathscr{P}(0,1)$ to avoid dealing with restricted systems of three $2$-sets with a common element.
-4.2.1 $\mathscr{P}(0,1)$ – There exists a pair of elements that appears in at least three distinct $3$-sets
+However, we prioritize the application of induction based on $\mathscr{P}(0,1)$ to avoid dealing with restricted CNFs of three $2$-clauses with a common variable.
+6.2.1 $\mathscr{P}(0,1)$ – There exists a pair of variables that appears in at least three distinct $3$-clauses
 Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,1)$.
-Let the distinct elements $a$ and $b$ appear together in the sets $\{a,b,c\},\{a,b,d\}$ and $\{a,b,e\}$ where $c$, $d$, and $e$ are distinct elements.
-Either one of $\{a,b\}$ is such that it always appears in a set together with the other element in $\{a,b\}$ or not.
+Let the distinct variables $a$ and $b$ appear together in the clauses $\{a,b,c\},\{a,b,d\}$ and $\{a,b,e\}$ where $c$, $d$, and $e$ are distinct variables.
+Either one of $\{a,b\}$ is such that it always appears in a clause together with the other variable in $\{a,b\}$ or not.
 We consider each case in turn and work with the corresponding anchors.
@@
 Without loss of generality (due to symmetry between $a$ and $b$), assume that $a$ always appears together with $b$.
-Core elements Forced elements Subsystem parameters Inductive bound
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total $\frac{31}{36}$ $\frac{79}{84}$
-4.2.1.2 $a$ and $b$ each appear in a set without the presence of the other element.
-We now extend the anchor by including sets containing either $a$ or $b$ without the other element being present.
-Observe that when $a=0$ and $b=1$ in the following table, we guaranteed the existence of at least one $2$-set in the resulting restricted system.
-Core elements Forced elements Subsystem parameters Inductive bound
+6.2.1.2 $a$ and $b$ each appear in a clause without the presence of the other variable.
+We now extend the anchor by including clauses containing either $a$ or $b$ without the other variable being present.
+Observe that when $a=0$ and $b=1$ in the following table, we guaranteed the existence of at least one $2$-clause in the resulting restricted CNF.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total $\frac{11}{12}$ $\frac{79}{84}$
-4.2.2 $\mathscr{P}(0,2)$ – There exists an element that appears in at least three distinct $3$-sets
+6.2.2 $\mathscr{P}(0,2)$ – There exists a variable that appears in at least three distinct $3$-clauses
 Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,2)$.
-Let $a$ be an element that appears in three sets $\{a,b,c\}$, $\{a,d,e\}$ and $\{a,f,g\}$ of $\mathbb{F}$ where $\{b,c\}$, $\{d,e\}$ and $\{f,g\}$ are distinct and do not have a common element.
-Core elements Forced elements Subsystem parameters Inductive bound
+Let $a$ be an variable that appears in three clauses $\{a,b,c\}$, $\{a,d,e\}$ and $\{a,f,g\}$ of $\mathbb{F}$ where $\{b,c\}$, $\{d,e\}$ and $\{f,g\}$ are distinct and do not have a common variable.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total 11 11
-4.2.3 $\mathscr{P}(0,3)$ – There exists a pair of elements that appears in two distinct $3$-sets
+6.2.3 $\mathscr{P}(0,3)$ – There exists a pair of variables that appears in two distinct $3$-clauses
 Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,3)$.
-Assume that $\mathbb{F}$ has a pair $\{a,b\}$ of elements which appears in exactly two sets $\{a,b,c\}$ and $\{a,b,d\}$ since otherwise if $\{a,b\}$ were to appear in three sets, we would have applied induction using $\mathscr{P}(0,1)$.
-We also know that $a$ (and $b$) cannot appear in any other set since otherwise we would have applied induction using $\mathscr{P}(0,2)$.
-Core elements Forced elements Subsystem parameters Inductive bound
+Assume that $\mathbb{F}$ has a pair $\{a,b\}$ of variables which appears in exactly two clauses $\{a,b,c\}$ and $\{a,b,d\}$ since otherwise if $\{a,b\}$ were to appear in three clauses, we would have applied induction using $\mathscr{P}(0,1)$.
+We also know that $a$ (and $b$) cannot appear in any other clause since otherwise we would have applied induction using $\mathscr{P}(0,2)$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,4)$.
-Let $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ be a triangle configuration in $\mathbb{F}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
-Moreover, we assume that $a$, $b$, and $c$ do not appear in any other set of $\mathbb{F}$.
-Core elements Forced elements Subsystem parameters Inductive bound
+Let $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ be a triangle configuration in $\mathbb{F}$ where the clause $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
+Moreover, we assume that $a$, $b$, and $c$ do not appear in any other clause of $\mathbb{F}$.
+It is not the case that $d=e=f$ since $\mathscr{P}(0,3)$ is not applicable.
+However, two of the three variables $\{d,e,f\}$ can be the same.
+For the last three restrictions in the following table, $n^{\prime}$ could be n5n-5.
+However, due to monotonicity with respect to $n$, we can replace n5n-5 with n4n-4 without loss of generality.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 0 11 0 $f=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
-0 11 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
+0 0 11 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
 11 11 $c=e=f=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
-0 11 11 $a=d=f=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
+11 11 $a=d=f=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
 11 11 $b=e=d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{6}$ $\frac{1}{6}$
@@
 Let $\{a,b,c\}$ be the spine of the $\mathbf{\exists}$} configuration, and $\{a,d,e\},\{b,f,g\}$ and $\{c,h,i\}$ are its arms where the set $\{d,e,f,g,h,i\}$ is disjoint from $\{a,b,c\}$.
-We assume $a,b$, and $c$ do not appear in any other sets.
-Core elements Forced elements Subsystem parameters Inductive bound
+We assume $a,b$, and $c$ do not appear in any other clauses.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total $\frac{103}{108}$ $\frac{187}{189}$
-4.2.6 $\mathscr{P}(0,6)$ – There exists an element that appears only in one set
+6.2.6 $\mathscr{P}(0,6)$ – There exists a variable that appears only in one clause
 Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,6)$.
-Let $a$ be such that $a$ appears exactly once in $\mathbb{F}$ in the set $\{a,b,c\}$.
-We consider two cases: Each of $b$ and $c$ appear in a different set or one of them does not appear in any other set.
-4.2.6.1 Each of $b$ and $c$ appear in a different set:
+Let $a$ be such that $a$ appears exactly once in $\mathbb{F}$ in the clause $\{a,b,c\}$.
+We consider two cases: Each of $b$ and $c$ appear in a different clause or one of them does not appear in any other clause.
+6.2.6.1 Each of $b$ and $c$ appear in a different clause:
 Assume that $\mathbb{F}$ contains $\{a,b,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $d,e,f$ and $g$ are distinct and $a$ does not appear anywhere else.
-Core elements Forced elements Subsystem parameters Inductive bound
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total $\frac{283}{324}$ $\frac{19}{21}$
-4.2.6.2 One of $\{b,c\}$ does not appear in any other set:
-Assume that $\mathbb{F}$ contains $\{a,b,c\}$ and $a$ and $b$ do not appear anywhere else. Either $c$ does not appear in any other set or there is a set of the form $\{c,d,e\}$.
-Core elements Forced elements Subsystem parameters Inductive bound
+6.2.6.2 One of $\{b,c\}$ does not appear in any other clause:
+Assume that $\mathbb{F}$ contains $\{a,b,c\}$ and $a$ and $b$ do not appear anywhere else. Either $c$ does not appear in any other clause or there is a clause of the form $\{c,d,e\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
 $c$ $a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
 Total 11 11
-When $c=0$, the number of transversals when $c$ appears in another set is dominated by the number of transversals when $c$ does not appear in any other set.
+When $c=0$, the number of transversals when $c$ appears in another clause is dominated by the number of transversals when $c$ does not appear in any other clause.
 6.3 Proof of the Upper bounds for $\Phi_{1}(s,t)$
-Let $\Xi$} the family of $3$-set systems that contain a $2$-set.
-Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n$ elements such that $\mathbb{F}=\{a,b\}\cup\mathbb{F}^{\prime}$ with threshold $t$ and deficit $s$.
-The following properties determine the subclasses.
+We cover $\mathbb{F}\in{\Xi}_{1}(s,t)$ with the following properties.
 $\mathscr{P}(1,1)$:
@@
 $\mathscr{P}(1,2)$:
-At least one of the elements in $\{a,b\}$ appears at least twice.
-4.3.1 At least one variable in $\{a,b\}$ appears exactly once  – $\mathscr{P}(1,1)$
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-11 0 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
-0 11 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
+At least one of the variables in $\{a,b\}$ appears at least twice.
+6.3.1 $\mathscr{P}(1,1)$ – At least one variable in $\{a,b\}$ appears exactly once
+Assume $\mathbb{F}\in{\Xi}_{1}(s,t)$ satisfies $\mathscr{P}(1,1)$. Without loss of generality, assume that the variable $a$ appears only in the clause $\{a,b\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
 Total $\frac{14}{15}$ 11
-4.3.2 At least one of the elements in $\{a,b\}$ appears at least twice  – $(\neg\mathscr{P}(1,1))\land\mathscr{P}(1,2)$
-In this case, $\mathbb{F}$ has a set of the form $\{a,c,d\}$ where $\{c,d\}$ is disjoint from $\{a,b\}$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+6.3.2 $\mathscr{P}(1,2)$ – At least one of the variables in $\{a,b\}$ appears at least twice
+Assume $\mathbb{F}\in{\Xi}_{1}(s,t)$ satisfies $\mathscr{P}(1,1)$. In this case, $\mathbb{F}$ has a clause of the form $\{a,c,d\}$ where $\{c,d\}$ is disjoint from $\{a,b\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 n1n-1 t1t-1 s2s-2 0 $\frac{3}{5}$ $\frac{7}{12}$
-0 11 n2n-2 t1t-1 s1s-1 11 $\frac{2}{5}$ $\frac{5}{12}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 11 $\frac{2}{5}$ $\frac{5}{12}$
 Total 11 11
 6.4 Proof of the Upper bounds for $\Phi_{2o}(s,t)$
-We now prove the upper bounds for $3$-set systems which contain two overlapping $2$-sets as stated in Theorem 4.1.
-Let $\Xi$} be the class of $3$-set system with two overlapping $2$-sets.
-Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$.
-Without loss of generality, assume $\mathbb{F}$ contains the overlapping $2$-sets $\{a,b\}$ and $\{a,c\}$.
-We first deal with the simpler case of odd $s$. We will show that the number of $t$-transversals is at most $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Odd $s$
-11 n1n-1 t1t-1 s2s-2 $\frac{7}{10}$
-0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{3}{10}$
+We now prove the upper bounds for monotone $3$-CNFs which contain two overlapping $2$-clauses as stated in Theorem 4.1.
+Let $\mathbb{F}\in{\Xi}_{2o}(s,t)$ be a monotone $3$-CNF over $n\geq 3$ variables with threshold $t$ and deficit $s$.
+Without loss of generality, assume $\mathbb{F}$ contains the overlapping $2$-clauses $\{a,b\}$ and $\{a,c\}$.
+We will first assume $s$ is odd and show that the number of $t$-transversals is at most $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
+11 n1n-1 t1t-1 s2s-2 0 $\frac{7}{10}$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{3}{10}$
 Total 11
@@
 We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{2}{3})(\frac{2}{3})^{\frac{s}{2}}3^{t}$.
-For this purpose, we consider the following subclasses of $\Xi$}.
+For this purpose, we cover ${\Xi}_{2o}(s,t)$ with the following properties.
+Observe that if at least one of $b$ or $c$ appears at least three times, then one of $\mathscr{P}(2o,1),\ldots,\mathscr{P}(2o,8)$ holds.
+The remaining two properties cover the case when both $b$ and $c$ appear at most twice.
 $\mathscr{P}(2o,1)$:
@@
 $\mathscr{P}(2o,4)$:
-There are two sets containing both $b$ and $c$.
+There are two clauses containing both $b$ and $c$.
 $\mathscr{P}(2o,5)$:
-There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
+There is exactly one clause in which both $b$ and $c$ appear together and exactly two other clauses, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
 $\mathscr{P}(2o,6)$:
-There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
+There exist clauses $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another clause that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
 $\mathscr{P}(2o,7)$:
-For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
+For each $x\in\{b,c\}$, there are exactly two clauses other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
 $\mathscr{P}(2o,8)$:
-One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
+One of $b$ or $c$ appears exactly in three clauses which do not contain the other variable and the other variable appears exactly in two clauses which do not contain the first variable.
 $\mathscr{P}(2o,9)$:
-There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$.
+There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and exactly one clause that contains both $b$ and $c$.
 $\mathscr{P}(2o,10)$:
-There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together.
-Lemma 4.6.
-If $\mathbb{F}$ satisfies $\neg(\bigwedge_{j=1}^{3}\mathscr{P}(2o,j))$, then $a$ appears exactly twice and $b$ and $c$ appear at least two times and at most three times in $\mathbb{F}$.
-Moreover, the following scenarios cover all the situations if one of $b$ or $c$ were to appear three times.
-  1. 1.
-    There are two sets containing both $b$ and $c$.
-  2. 2.
-    There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
-  3. 3.
-    There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
-  4. 4.
-    For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
-  5. 5.
-    One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
-Lemma 4.7.
-If $\mathbb{F}$ satisfies $\neg\bigwedge_{j=1}^{8}\mathscr{P}(2o,j)$, then $a$, $b$ and $c$ appear exactly twice.
+There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and $b$ and $c$ do not appear together.
 6.4.1 $a$ appears three times  – $\mathscr{P}(2o,1)$
-Without of loss of generality, assume that $a$ appears in $\{a,d,e\}$ in addition to appearing in $\{a,b\}$ and $\{a,c\}$ where neither $d$ nor $e$ are in $\{b,c\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
-11 n1n-1 t1t-1 s2s-2 $\frac{3}{4}$
-0 $b=c=1$ n3n-3 t2t-2 s3s-3 1 $\frac{1}{4}$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,1)$.
+Without loss of generality, assume that $a$ appears in $\{a,d,e\}$ in addition to appearing in $\{a,b\}$ and $\{a,c\}$ where neither $d$ nor $e$ are in $\{b,c\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 n1n-1 t1t-1 s2s-2 0 $\frac{3}{4}$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
 Total 11
-4.4.2 Either $b$ or $c$ appears four times  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2o,j))\land\mathscr{P}(20,2)$
+6.4.2 $\mathscr{P}(2o,2)$ – Either $b$ or $c$ appears four times
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,2)$.
 Without of loss of generality, assume that $b$ appears in $\{a,b\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{b,h,i\}$ where none of $\{d,e\}$, $\{f,g\}$ and $\{h,i\}$ is equal to $\{a,c\}$.
@@
 6.4.2.1 Case 1:
-There is no element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+There is no variable common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 n1n-1 t1t-1 s2s-2 11 $\frac{5}{8}$
@@
 6.4.2.2 Case 2:
-There is an element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$. Let $d=f=h$
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
+There is a variable common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$. Let $d=f=h$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$b$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 n1n-1 t1t-1 s2s-2 11 $\frac{5}{8}$
-0 0 $a=d=f=h=1$ n6n-6 t4t-4 s6s-6 0 $\frac{1}{16}$
+0 0 $a=e=g=i=1$ n6n-6 t4t-4 s6s-6 0 $\frac{1}{16}$
 0 11 $a=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
 Total $\frac{47}{48}$
-4.4.3 One of $b$ or $c$ appears only once  – $(\neg\bigvee_{j=1}^{2}\mathscr{P}(2o,j))\land\mathscr{P}(2o,3)$
-Assume that $b$ appears only once.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
-11 0 n2n-2 t1t-1 s1s-1 $\frac{7}{12}$
-0 11 $c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
+6.4.3 $\mathscr{P}(2o,3)$ – One of $b$ or $c$ appears only once
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,3)$. Assume without loss of generality that $b$ appears only once.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 0 n2n-2 t1t-1 s1s-1 0 $\frac{7}{12}$
+0 11 $c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
 Total $\frac{7}{8}$
-4.4.4 There are two sets containing both $b$ and $c$  – $(\neg\bigvee_{j=1}^{3}\mathscr{P}(2o,j))\land\mathscr{P}(2o,4)$
-Without of loss of generality, assume we have two sets $\{b,c,d\}$ and $\{b,c,e\}$ where $d\neq e$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
-11 0 0 $d=e=1$ n5n-5 t3t-3 s4s-4 $\frac{1}{8}$
-11 0 11 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
-11 11 0 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
-0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
+6.4.4 $\mathscr{P}(2o,4)$ – There are two clauses containing both $b$ and $c$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,4)$. Without of loss of generality, assume we have two clauses $\{b,c,d\}$ and $\{b,c,e\}$ where $d\neq e$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 0 0 $d=e=1$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{8}$
+11 0 11 n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
+11 11 0 n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
 Total 11
-4.4.5 There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$  – $(\neg\bigvee_{j=1}^{4}\mathscr{P}(2o,j))\land\mathscr{P}(2o,5)$
-Without of loss of generality, assume we have a set $\{b,c,d\}$.
-Furthermore, assume that there is another set $\{b,e,f\}$ where neither $e$ nor $f$ is equal to $c$ and a set $\{c,g,h\}$ where neither $g$ nor $h$ is equal to $b$.
+6.4.5 $\mathscr{P}(2o,5)$ – There is exactly one clause in which both $b$ and $c$ appear together and exactly two other clauses, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,5)$.
+Without of loss of generality, assume we have a clause $\{b,c,d\}$.
+Furthermore, assume that there is another clause $\{b,e,f\}$ where neither $e$ nor $f$ is equal to $c$ and a clause $\{c,g,h\}$ where neither $g$ nor $h$ is equal to $b$.
 We consider two cases:
-4.4.5.1 $d\not\in\{e,f,g,h\}$:
-$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+6.4.5.1 Case 1: $d\not\in\{e,f,g,h\}$:
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 $a=1$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
@@
 Total 11
-4.4.5.2 $d\in\{e,f,g,h\}$:
-Assume that $d=e$. In this case, we have the sets $\{a,b\},\{a,c\},\{b,c,d\}$, $\{b,d,f\}$ and $\{c,g,h\}$.
-$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+6.4.5.2 Case 2: $d\in\{e,f,g,h\}$:
+Assume that $d=e$. In this case, we have the clauses $\{a,b\},\{a,c\},\{b,c,d\}$, $\{b,d,f\}$ and $\{c,g,h\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 $a=1,d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
@@
 Total $\frac{11}{12}$
-4.4.6 There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$   – $(\neg\bigvee_{j=1}^{5}\mathscr{P}(2o,j))\land\mathscr{P}(2o,6)$
-Assume that there are sets $\{b,d,e\}$, $\{c,d,e\}$, and $\{b,f,g\}$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
-11 0 n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
+6.4.6 $\mathscr{P}(2o,6)$ – There exist clauses $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another clause that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,6)$. Assume that there are clauses $\{b,d,e\}$, $\{c,d,e\}$, and $\{b,f,g\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 0 n2n-2 t1t-1 s1s-1 22 $\frac{5}{12}$
 11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
@@
 Total $\frac{23}{24}$
-4.4.7 For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$  – $(\neg\bigvee_{j=1}^{6}\mathscr{P}(2o,j))\land\mathscr{P}(2o,7)$
+6.4.7 $\mathscr{P}(2o,7)$ – For each $x\in\{b,c\}$, there are exactly two clauses other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,7)$.
 Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$, $\{c,d^{\prime},e^{\prime}\}$ and $\{c,f^{\prime},g^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$ are distinct.
 We consider three cases:
-4.4.7.1 Case 1:
-There is no element common to three or more of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+6.4.7.1 Case 1: There is no variable common to three or more of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 0 n3n-3 t1t-1 $s$ 33 $\frac{7}{24}$
-11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
-11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
+11 0 11 n3n-3 t3t-3 s3s-3 22 $\frac{5}{24}$
+11 11 0 n3n-3 t3t-3 s3s-3 22 $\frac{5}{24}$
 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
 Total 11
-4.4.7.2 Case 2:
-There is an element common to exactly three of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
+6.4.7.2 Case 2: There is a variable common to exactly three of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
 Assume without loss of generality, $d=f=d^{\prime}$ where $d$ is distinct from $f^{\prime}$ and $g^{\prime}$.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
 11 0 0 11 n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
-11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
-11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
+11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{24}$
+11 11 0 n3n-3 t3t-3 s3s-3 22 $\frac{5}{24}$
 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
 Total 11
-4.4.7.3 Case 3:
-There is an element common to exactly four of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$. Assume without loss of generality, $d=f=d^{\prime}=f^{\prime}$.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
-11 0 0 0 $e=g=e^{\prime}=g^{\prime}=1$ n8n-8 t5t-5 s7s-7 0 $\frac{1}{24}$
-11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
-11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
-11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
+6.4.7.3 Case 3: There is a variable common to exactly four of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
+Assume without loss of generality, $d=f=d^{\prime}=f^{\prime}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 0 0 0 $e=g=e^{\prime}=g^{\prime}=1$ n8n-8 t5t-5 s7s-7 0 $\frac{7}{288}$
+11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
+11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{24}$
+11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{24}$
 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
-Total 11
-4.4.8 One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element  – $(\neg\bigvee_{j=1}^{7}\mathscr{P}(2o,j))\land\mathscr{P}(2o,8)$
+Total $\frac{283}{288}$
+6.4.8 $\mathscr{P}(2o,8)$ – One of $b$ or $c$ appears exactly in three clauses which do not contain the other variable and the other variable appears exactly in two clauses which do not contain the first variable
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,8)$.
 Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{c,d^{\prime},e^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$ are distinct.
 We consider two cases:
-Case 1: There is no element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+6.4.8.1 Case 1: There is no variable that is common to the clauses, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 0 n3n-3 t1t-1 $s$ 33 $\frac{7}{24}$
+11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 22 $\frac{5}{36}$
+11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
+Total $\frac{35}{36}$
+6.4.8.2 Case 2: There is a variable that is common to the clauses, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
+Assume $d=f=d^{\prime}$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
+11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
 11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 $2o$ $\frac{5}{36}$
@@
 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
-Total $\frac{35}{36}$
-Case 2: There is an element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$. Assume $d=f=d^{\prime}$
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
-11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
-11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
-11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 $2o$ $\frac{5}{36}$
-11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
-0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
-Total $\frac{35}{36}$
-4.4.9 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$  – $(\neg\bigvee_{j=1}^{8}\mathscr{P}(2o,j))\land\mathscr{P}(2o,9)$
-Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\}$, and $\{b,c,d\}$ where $d$ is neither $b$ nor $c$. We consider two cases.
-4.4.9.1 $d$ appears uniquely:
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+Total $\frac{47}{48}$
+6.4.9 $\mathscr{P}(2o,9)$ – There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and exactly one clause that contains both $b$ and $c$
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,9)$. Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\}$, and $\{b,c,d\}$ where $d$ is neither $b$ nor $c$. We consider two cases.
+6.4.9.1 Case 1: $d$ appears uniquely:
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n$ $t$ $s$ Type Even $s$
 11 0 0 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
@@
 Total 11
-4.4.9.2 $d$ does not appear uniquely:
+6.4.9.2 Case 2: $d$ does not appear uniquely:
 There is another clause $\{d,f,g\}$ where $f$ and $g$ are distinct from $b$ and $c$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n$ $t$ $s$ Type Even $s$
 11 0 0 $d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
 11 0 11 $d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
-11 11 $c=d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{1}{4}$
+11 11 $c=d=0$ n4n-4 t2t-2 s2s-2 11 $\frac{5}{24}$
 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
-Total 11
-4.4.10 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together  – $(\neg\bigvee_{j=1}^{9}\mathscr{P}(2o,j))\land\mathscr{P}(2o,10)$
-Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $\{d,e,f,g\}\cap\{b,c\}=\varnothing$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
+Total $\frac{23}{24}$
+6.4.10 $\mathscr{P}(2o,10)$ – There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and $b$ and $c$ do not appear together
+Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,10)$.
+Assume without loss of generality that $b$ and $c$ only appear in $\{a,b\},\{a,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $\{d,e,f,g\}\cap\{b,c\}=\varnothing$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 11 0 n2n-2 t1t-1 s1s-1 11 $\frac{1}{2}$
@@
 6.5 Proof of the Upper bounds for $\Phi_{2d}(s,t)$
-We now prove the upper bounds for $3$-set systems which contain two disjoint $2$-sets as stated in Theorem 4.1.
-Let $\Xi$} be the class of $3$-set system with two disjoint $2$-sets.
-Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 4$ elements with threshold $t$ and deficit $s$.
-Without loss of generality, assume $\mathbb{F}$ contains two disjoint $2$-sets $\{a,b\}$ and $\{c,d\}$.
+We now prove the upper bounds for monotone $3$-CNFs which contain two disjoint $2$-clauses as stated in Theorem 4.1.
+Let ${\Xi}_{2d}(s,t)$ be the class of monotone $3$-CNFs over $n\geq 4$ variables with threshold $t$ and deficit $s$.
+Without loss of generality, assume $\mathbb{F}$ contains two disjoint $2$-clauses $\{a,b\}$ and $\{c,d\}$.
 We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{5}{6})^{2}(\frac{2}{3})^{\frac{s}{2}}3^{t}$ when $s$ is even and by $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$ when $s$ is odd.
-We cover $\Xi$} based on the following properties.
+We cover ${\Xi}_{2d}(s,t)$ with the following properties.
 $\mathscr{P}(2d,1)$:
@@
 $\mathscr{P}(2d,2)$:
-There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$.
+There is a clause containing a pair of variables from $\{a,b\}\times\{c,d\}$.
 $\mathscr{P}(2d,3)$:
-A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$.
+A pair of variables appears in two clauses where the third variable is from $\{a,b,c,d\}$.
 $\mathscr{P}(2d,4)$:
@@
 $\mathscr{P}(2d,5)$:
-There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
+There exists two clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
 Lemma 6.8.
-If $\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j)$, then $a$ and $b$ must each appear in another set which contains neither $c$ not $d$ and the two sets cannot have more than one common element.
+If $\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j)$, then $a$ and $b$ must each appear in another clause which contains neither $c$ nor $d$ and the two clauses cannot have more than one common variable.
 Proof.
@@
   3. 3.
-    no pair from $\{a,b\}\times\{c,d\}$ appears in any set.
-Since $a$ and $b$ must participate in two sets, let there be two sets $\{a,e,f\}$ and $\{b,g,h\}$ which are distinct from $\{a,b\}$ and $\{c,d\}$.
-It must be $\{e,f,g,h\}\cap\{a,b,c,d\}=\varnothing$ since no pair from $\{a,b\}\times\{c,d\}$ appears in any set.
+    no pair from $\{a,b\}\times\{c,d\}$ appears in any clause.
+Since $a$ and $b$ must participate in two clauses, let there be two clauses $\{a,e,f\}$ and $\{b,g,h\}$ which are distinct from $\{a,b\}$ and $\{c,d\}$.
+It must be $\{e,f,g,h\}\cap\{a,b,c,d\}=\varnothing$ since no pair from $\{a,b\}\times\{c,d\}$ appears in any clause.
 Moreover, $\{e,f\}\neq\{g,h\}$ since no pair occurs with one of $\{a,b,c,d\}$ more than once.
 ∎
-4.5.1 One of $\{a,b,c,d\}$ appears 3 times  – $\mathscr{P}(2d,1)$
-Assume that $\mathbb{F}$ contains an element in $\{a,b,c,d\}$, say $a$, which appears in three sets.
-Let $\{a,e,f\}$ and $\{a,g,h\}$ be the other two sets in which $a$ appears.
+6.5.1 $\mathscr{P}(2d,1)$ – One of $\{a,b,c,d\}$ appears 3 times
+Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,1)$.
+Assume that $\mathbb{F}$ contains an variable in $\{a,b,c,d\}$, say $a$, which appears in three clauses.
+Let $\{a,e,f\}$ and $\{a,g,h\}$ be the other two clauses in which $a$ appears.
 Neither $\{e,f\}$ nor $\{g,h\}$ is equal to $\{c,d\}$.
-We consider two cases.
-4.5.1.1 Case 1:
-There is no common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+We consider two cases
+6.5.1.1 Case 1: There is no common variable for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
@@
 Total $\frac{24}{25}$ $\frac{19}{20}$
-4.5.1.2 Case 2:
-There is a common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$. Let $e=f=d$.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+6.5.1.2 Case 2: There is a common variable for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
+Let $e=f=d$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
@@
 Total $\frac{47}{50}$ $\frac{23}{24}$
-4.5.2 There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2d,j))\land\mathscr{P}(2d,2)$
-   Assume without loss of generality that $\mathbb{F}$ contains a set of the form $\{a,c,e\}$ where $e$ is distinct from $b$ and $d$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+6.5.2 $\mathscr{P}(2d,2)$ – There is a clause containing a pair of variables from $\{a,b\}\times\{c,d\}$
+   Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,2)$.
+Assume without loss of generality that $\mathbb{F}$ contains a clause of the form $\{a,c,e\}$ where $e$ is distinct from $b$ and $d$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 n1n-1 t1t-1 s2s-2 11 $\frac{3}{5}$ $\frac{3}{5}$
-0 1 n2n-2 t1t-1 s1s-1 22 overlapping $\frac{2}{5}$ $\frac{2}{5}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{2}{5}$ $\frac{2}{5}$
 Total 11 11
-4.5.3 A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$  – $(\neg\bigwedge_{j=1}^{2}\mathscr{P}(2d,j))\land\mathscr{P}(2d,3)$
-Assume that there are two sets $\{e,f,g\}$ and $\{e,f,h\}$ containing the pair $\{e,f\}$ where $g,h\in\{a,b,c,d\}$.
+6.5.3 $\mathscr{P}(2d,3)$ – A pair of variables appears in two clauses where the third variable is from $\{a,b,c,d\}$
+Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,3)$.
+Assume that there are two clauses $\{e,f,g\}$ and $\{e,f,h\}$ containing the pair $\{e,f\}$ where $g,h\in\{a,b,c,d\}$.
 Since each of $\{a,b,c,d\}$ appear at most twice, we know that neither $e$ nor $f$ are in $\{a,b,c,d\}$.
-There are two cases.
-4.5.3.1 $g=a$ and $h=c$:
-$a$ $c$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+There are three cases.
+6.5.3.1 Case 1: $g=a$ and $h=c$:
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 0 $d=1,b=0$ n4n-4 t2t-2 s2s-2 1 $\frac{1}{5}$ $\frac{1}{5}$
@@
 Total $\frac{24}{25}$ $\frac{19}{20}$
-4.5.3.2 $g=a$ and $h=b$:
-$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+6.5.3.2 Case 2: $g=a$ and $h=b$:
+Further assume that one of $\{e,f\}$ does not appear in any other clause.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $e$ $f$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 0 0 11 n4n-4 t2t-2 s2s-2 11 $\frac{1}{5}$ $\frac{1}{5}$
@@
 Total 11 11
-4.5.4 One of $\{a,b,c,d\}$ appears exactly once  – $(\neg\bigwedge_{j=1}^{3}\mathscr{P}(2d,j))\land\mathscr{P}(2d,4)$
-Assume that $a$ is unique.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
+6.5.3.3 Case 3: $g=a$ and $h=b$:
+Further assume that both $e$ and $f$ appear in some other clause.
+Assume that there are clauses $\{e,u,v\}$ and $\{f,x,y\}$.
+Since we also have that the pair $\{e,f\}$ cannot appear in any other clause, it must be $\{u,v,x,y\}$ is disjoint from $\{e,f\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $e$ $f$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+11 0 11 n3n-3 t2t-2 s3s-3 11 $\frac{6}{25}$ $\frac{1}{4}$
+11 0 0 1 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
+0 11 11 n3n-3 t2t-2 s3s-3 11 $\frac{6}{25}$ $\frac{1}{4}$
+0 11 0 1 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
+11 11 0 0 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
+Total $\frac{49}{50}$ 11
+6.5.4 $\mathscr{P}(2d,4)$ – One of $\{a,b,c,d\}$ appears exactly once
+Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,4)$. Assume that $a$ is unique.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
 11 0 n2n-2 t1t-1 s1s-1 11 $\frac{12}{25}$ $\frac{1}{2}$
@@
 Total $\frac{24}{25}$ 11
-4.5.5 There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$  – $(\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j))\land\mathscr{P}(2d,5)$
-We have that every one of $\{a,b,c,d\}$ appears exactly twice in $\mathbb{F}$. Assume we have sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
-$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-11 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
-0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
+6.5.5 $\mathscr{P}(2d,5)$ – There exists two clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$
+Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,5)$.
+We have that every one of $\{a,b,c,d\}$ appears exactly twice in $\mathbb{F}$.
+Assume we have clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$ (that is, at least one of $e,f$ is neither $g$ nor $h$, but $e,f,g,h$ may be three total variables).
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+11 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{2}{5}$ $\frac{5}{12}$
+0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{2}{5}$ $\frac{5}{12}$
 11 1 $e=f=g=h=0$ n5n-5 t2t-2 s1s-1 11 $\frac{4}{25}$ $\frac{1}{6}$
@@
 6.6 Proof of the Upper bounds for $\Phi_{3}(s,t)$
-We now prove the upper bounds for $3$-set systems which contain three $2$-sets as stated in Theorem 4.1.
-Let $\Xi$} be the class of $3$-set system with three $2$-sets.
-Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$.
-Without loss of generality, assume $\mathbb{F}$ contains the $2$-sets $\{a,b\}$, $\{c,d\}$, and $\{e,f\}$ where $a,b,c,d,e$ and $f$ are not necessarily distinct.
-We regard the $2$-sets as edges of a graph, consider several cases based on the structure of the graph and prove the bounds in each case.
-4.6.1 The three 2-sets form a path of length 3
-Consider the case where $F$ contains three $2$-sets that form a path of length $3$.
-Let $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$ be the $2$-sets that form a path (of length $3$).
-We cover the family of such set systems by the following sub families.
-$\mathscr{P}(3,1)$:
-At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
-$\mathscr{P}(3,2)$:
-Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$
-Fact 4.9.
-Either $\mathscr{P}(3,1)$ or $\mathscr{P}(3,2)$ holds.
-4.6.1.1 At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.  – $\mathscr{P}(3,1)$
-Assume without loss of generality that $b$ appears in no other set.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
-0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
-1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
+We now prove the upper bounds for monotone $3$-CNFs which contain three $2$-clausess as stated in Theorem 4.1.
+Let $\mathbb{F}\in{\Xi}_{3}(s,t)$ be the class of such monotone $3$-CNFs over $n\geq 3$ variables with threshold $t$ and deficit $s$.
+Without loss of generality, assume $\mathbb{F}$ contains the $2$-clauses $\{a,b\}$, $\{c,d\}$, and $\{e,f\}$ where $a,b,c,d,e$ and $f$ are not necessarily distinct.
+We regard the $2$-clauses as edges of a graph, consider several cases based on the structure of the graph and prove the bounds in each case.
+6.6.1 The three $2$-clauses form a path of length 3
+Consider the case where $\mathbb{F}$ contains three $2$-clauses that form a path of length $3$.
+Let $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$ be the $2$-clauses that form a path (of length $3$).
+We cover the family of such monotone $3$-CNFs with the following sub families.
+$\mathscr{P}(3\text{-}{p},1)$:
+At least one of $\{b,c\}$ is in no other clause than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
+$\mathscr{P}(3\text{-}{p},2)$:
+Both $b$ and $c$ appear in some other clause other than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$
+6.6.1.1 $\mathscr{P}(3\text{-}{p},1)$ – At least one of $\{b,c\}$ is in no other clause than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
+Assume without loss of generality that $b$ does not appear in any other clause.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+1 $b=0,c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
+0 1 $b=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
+0 0 $b=d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{2}{7}$ $\frac{7}{27}$
+Total $\frac{20}{21}$ $\frac{25}{27}$
+6.6.1.2 $\mathscr{P}(3\text{-}{p},2)$ – Both $b$ and $c$ appear in some other clause other than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
+Assume without loss of generality that $b$ appears in another clause which is distinct from the clauses in the path. Let $\{b,e,f\}$ such a clause. Note that $e$ and $f$ are distinct from $a$ and $c$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+1 n1n-1 t1t-1 s2s-2 11 $\frac{5}{7}$ $\frac{2}{3}$
+0 $a=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
+Total 11 $\frac{17}{18}$
+6.6.2 The three $2$-clauses form a triangle
+Consider the case where $\mathbb{F}\in{\Xi}_{3}(s,t)$ contains three $2$-clauses that form a triangle (cycle of length 3) configuration.
+Let $\{a,b\}$, $\{b,c\}$ and $\{c,a\}$ be the $2$-clauses that form form the triangle.
+We cover ${\Xi}_{3}(s,t)$ based on the following properties.
+$\mathscr{P}(3\text{-}{t},1)$:
+At least one of $\{a,b,c\}$ is is no other clause outside the triangle.
+$\mathscr{P}(3\text{-}{t},2)$:
+At least one of $\{a,b,c\}$ is in exactly one other clause outside the triangle, and all are in at least one other clause.
+$\mathscr{P}(3\text{-}{t},3)$:
+Each of $\{a,b,c\}$ are in at least two other clauses outside the triangle.
+6.6.2.1 $\mathscr{P}(3\text{-}{t},1)$ – At least one of $\{a,b,c\}$ is is no other clause outside the triangle
+We must select at least two variables of the triangle, otherwise either one of the clauses would be unsatisfied.
+Assume without loss of generality that $a$ is in no other clause outside of the triangle.
+In this case, not all the variables in the triangle can be clause to 1 simultaneously.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
+1 0 $c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
+1 1 $c=0$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
 Total 11 11
-4.6.1.2 Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$  – $\mathscr{P}(3,2)$
-Assume without loss of generality that $b$ appears in another set which is distinct from the sets in the path. Let $\{b,e,f\}$ such a set. Note the $e$ and $f$ are distinct from $a$ and $c$.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-1 $\{c,d\}$ remains n1n-1 t1t-1 s2s-2 11 $\frac{5}{7}$ $\frac{2}{3}$
-1 0 1 $\{e,f\}$ remains n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
-Total 11 $\frac{17}{18}$
-4.6.2 The three 2-sets form a triangle
-Let $\mathbb{F}$ be a set system with three 2-sets that form a triangle (cycle of length 3) configuration.
-Let $\Xi$} be the class of such 3-set systems.
-We cover $\Xi$} based on the following properties.
-$\mathscr{P}(3,1)$:
-At least one of $\{a,b,c\}$ is is no other set outside the triangle.
-$\mathscr{P}(3,2)$:
-At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set.
-$\mathscr{P}(3,3)$:
-Each of $\{a,b,c\}$ are in at least two other sets outside the triangle.
-Lemma 4.10.
-One of $\mathscr{P}(3,1),\mathscr{P}(3,2),\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
-Proof.
-In this case, by $\neg\mathscr{P}(3,1)$ $a,b,c$ are each in at least one clause besides the triangle.
-By $\neg\mathscr{P}(3,2)$, then, we know none of them is in exactly one clause besides the triangle, as all are in at least one.
-This means each is in at least two additional clauses, implying $\mathscr{P}(3,3)$.
-By contradiction, this therefore cannot occur.
-∎
-4.6.2.1 At least one of $\{a,b,c\}$ is is no other set outside the triangle  – $\mathscr{P}(3,1)$
-We must select at least two elements of the triangle, otherwise either one of the clauses would be unsatisfied.
-Assume without loss of generality that $a$ is in no other set outside of the triangle.
-In this case, not all the elements in the triangle can be set to 1 simultaneously.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
-1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
-1 1 0 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
-Total 11 11
-4.6.2.2 At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set  – $\mathscr{P}(3,2)$
-Assume without loss of generality that $a$ is in exactly in one set $\{a,d,e\}$ outside the triangle where $d$ and $e$ are distinct from $b$ and $c$. Note that this set must not contain $b$ or $c$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-0 1 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
-1 0 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
+6.6.2.2 $\mathscr{P}(3\text{-}{t},2)$ – At least one of $\{a,b,c\}$ is in exactly one other clause outside the triangle, and all are in at least one other clause
+Assume without loss of generality that $a$ is in exactly one clause $\{a,d,e\}$ outside the triangle where $d$ and $e$ are distinct from $b$ and $c$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
+1 0 $c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
 1 1 0 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
@@
 Total 11 $\frac{26}{27}$
-4.6.2.3 Each of $\{a,b,c\}$ are in at least two other sets outside the triangle  – $\mathscr{P}(3,3)$
+6.6.2.3 $\mathscr{P}(3\text{-}{t},3)$ – Each of $\{a,b,c\}$ are in at least two other clauses outside the triangle
 The two clauses attached to any triangle member must not be redundant with each other, and they cannot contain other triangle variables without being redundant.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
-0 1 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
-1 0 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
+0 $b=c=1$ n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
+1 0 $c=1$ n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
 1 1 0 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
@@
 Total $\frac{13}{14}$ $\frac{8}{9}$
-4.6.3 There exists a 2-set that is disjoint from the other two 2-sets
-Let $\mathbb{F}$ be a set system with three 2-sets that where one of the 2-sets is disjoint from the remaining 2-set.
-Without loss of generality, let this isolated 2-set be $\{a,b\}$.
-Let $\{c,d\}$ and $\{e,f\}$ be the other two sets.
-We will first deal with the cases when $s$ is odd. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
-$\mathscr{P}(3,1)$:
-$s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap.
-$\mathscr{P}(3,2)$:
-$s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$.
-$\mathscr{P}(3,3)$:
-$s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$.
+6.6.3 There exists a $2$-clause that is disjoint from the other two $2$-clauses
+Let $\mathbb{F}$ be a monotone $3$-CNF with three $2$-clauses that where one of the 2-clauses is disjoint from the remaining $2$-clauses.
+Let ${\Xi}_{3\text{-}i}(s,t)$ be the class of such CNFs.
+Without loss of generality, let this isolated $2$-clauses be $\{a,b\}$.
+Let $\{c,d\}$ and $\{e,f\}$ be the other two $2$-clauses.
+We will first deal with the cases when $s$ is odd. We cover such $\mathbb{F}$ with the following properties.
+$\mathscr{P}(3\text{-}{io},1)$:
+$s$ is odd and the other two clauses $\{c,d\}$ and $\{e,f\}$ overlap.
+$\mathscr{P}(3\text{-}{io},2)$:
+$s$ is odd, $a$ appears in no other clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
+$\mathscr{P}(3\text{-}{io},3)$:
+$s$ is odd, $a$ appears in some other 3-clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
+6.6.3.1 $\mathscr{P}(3\text{-}{io},1)$$s$ is odd and the other two clauses $\{c,d\}$ and $\{e,f\}$ overlap.
+If $\{c,d\}$ and $\{e,f\}$ overlap, the following calculation applies.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
+1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{5}{9}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{4}{9}$
+Total 11
+6.6.3.2 $\mathscr{P}(3\text{-}{io},2)$$s$ is odd, $a$ appears in no other clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
+We take advantage of the fact that $a$ appears uniquely since $a$ and $b$ cannot be simultaneously $1$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
+1 $b=0$ n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
+Total $\frac{25}{27}$
+6.6.3.3 $\mathscr{P}(3\text{-}{io},3)$$s$ is odd, $a$ appears in some other 3-clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
+In this case, no single added 2-clause can possibly make the three form a star, as the existing two don’t share any points of overlap. We can therefore use the standard type 3 bound.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
+1 n1n-1 t1t-1 s2s-2 $2d$ $\frac{5}{9}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
+Total $\frac{17}{18}$
+We now deal with the case that $s$ is even. We cover ${\Xi}_{3\text{-}i}(s,t)$ with the following properties when $s$ is even.
+$\mathscr{P}(3\text{-}{ie},1)$:
+$s$ is even and either $a$ or $b$ appear in no other clause.
+$\mathscr{P}(3\text{-}{ie},2)$:
+$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in at most one of $cd$ and $ef$.
+$\mathscr{P}(3\text{-}{ie},3)$:
+$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in both of $cd$ and $ef$.
+$\mathscr{P}(3\text{-}{ie},4)$:
+$s$ is even, $|\{c,d,e,f\}|=3$ and there exists clause $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
+$\mathscr{P}(3\text{-}{ie},5)$:
+$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist clauses $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
+$\mathscr{P}(3\text{-}{ie},6)$:
+$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist clauses $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
 Lemma 6.9.
-One of $\mathscr{P}(3,1),\mathscr{P}(3,2)$ and $\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
-4.6.3.1 $s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap  – $\mathscr{P}(3,1)$
-If $\{c,d\}$ and $\{e,f\}$ overlap, the following calculation applies.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
-1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{5}{9}$
-0 1 n2n-2 t1t-1 s1s-1 $2o$ $\frac{4}{9}$
+For $i\in[6]$, one of $\mathscr{P}(3\text{-}{ie},i)$ always holds for every monotone $3$-CNF in ${\Xi}_{3\text{-}i}(s,t)$ when $s$ is even.
+Proof.
+For even $s$, let $F\in{\Xi}_{3\text{-}i}(s,t)$ be arbitrary and assume none of $\mathscr{P}(3\text{-}{ie},i)$ holds for any choice of isolated $2$-clause in the three $2$-clause set.
+Since $\mathscr{P}(3\text{-}{ie},1)$ and $\mathscr{P}(3\text{-}{ie},2)$ don’t hold, $a$ and $b$ appear in at least one other clause, and all such clauses don’t contain any of $c,d,e,f$.
+Notice that among $c,d,e,f$, by non-redundancy, either there exist $3$ distinct variables or $4$.
+If there were $3$, then we see that $\mathscr{P}(3\text{-}{ie},4)$ must have been satisfied.
+Hence, it must be the case that $c,d,e,f$ are all distinct.
+This means that the three 2-sets in $F$ must be disjoint.
+Therefore by symmetry, instead of letting $\{a,b\}$ be the isolated 2-set, we can relabel variables and let any of $\{c,d\}$ or $\{e,f\}$ be the isolated 2-set and check for these properties.
+Hence, we can conclude that none of $a,b,c,d,e,f$ appears in any other set with each other (this is ruled out already for say $a,b$ by non-redundancy and remaining by negation of $\mathscr{P}(3\text{-}{ie},2)$ and symmetry).
+Also, we must have that each of $a,b,c,d,e,f$ must appear in at least one more 3-set.
+Since $\mathscr{P}(3\text{-}{ie},6)$ does not hold, it must be that both $a,b$ appear in exactly one other clause and by symmetry that each of $c,d,e,f$ appear in exactly one other clause.
+Also, since $\mathscr{P}(3\text{-}{ie},5)$ does not hold, we conclude that if the unique other 3-set where $a$ appears is say $\{a,x,y\}$ then it must be the case that there exists 3-set $\{c,x,y\}$.
+Symmetrically applying this argument to all pairs, we infer that case $\mathscr{P}(3\text{-}{ie},7)$ must hold for $F$, which is a contradiction.
+∎
+6.6.3.4 $\mathscr{P}(3\text{-}{ie},1)$$s$ is even and either $a$ or $b$ appear in no other clause.
+Without loss of generality let this be $a$. We take advantage of the fact that $a$ appears uniquely.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+1 $b=0$ n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
+Total $\frac{20}{21}$
+6.6.3.5 $\mathscr{P}(3\text{-}{ie},2)$$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in at most one of $cd$ and $ef$.
+Without loss of generality, let this be $a$ and $c$. Let this set be $\{a,c,x\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
+1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
+0 1 $b=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$
+0 0 $b=d=x=1$ n5n-5 t3t-3 s4s-4 11 $\frac{5}{42}$
 Total 11
-4.6.3.2 $s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,2)$
-We take advantage of the fact that $a$ appears uniquely since $a$ and $b$ cannot be simultaneously $1$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
-1 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
-0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
-Total $\frac{25}{27}$
-4.6.3.3 $s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
-In this case, no single added 2-clause can possibly make the three form a star, as the existing two don’t share any points of overlap. We can therefore use the standard type 3 bound.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
-1 n1n-1 t1t-1 s2s-2 $2d$ $\frac{5}{9}$
-0 1 n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
-Total $\frac{17}{18}$
-We now deal with the case that $s$ is even. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
-$\mathscr{P}(3,1)$:
-$s$ is even and either $a$ or $b$ appear in no other set.
-$\mathscr{P}(3,2)$:
-$s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$.
-$\mathscr{P}(3,3)$:
-$s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
-$\mathscr{P}(3,4)$:
-$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
-$\mathscr{P}(3,5)$:
-$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
-$\mathscr{P}(3,6)$:
-$s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists.
-Lemma 4.12.
-For $i\in[6]$, one of $\mathscr{P}(3,i)$ always holds for every set system in $\Xi$}.
-Proof.
-Let $F\in{\Xi}$ be arbitrary and assume none of $\mathscr{P}(3,i)$ holds for any choice of isolated 2-set in the three 2-set system.
-Since $\mathscr{P}(3,1)$ and $\mathscr{P}(3,2)$ don’t hold, $a$ and $b$ appear in at least one other clause, and all such clauses don’t contain any of $c,d,e,f$.
-Notice that amongst $c,d,e,f$, by non-redundancy, either there exist $3$ distinct variables or $4$.
-If there were $3$, then we see that $\mathscr{P}(3,3)$ must have been satisfied.
-Hence, it must be the case that $c,d,e,f$ are all distinct.
-This means that the three 3-sets in $F$ must be disjoint.
-Therefore by symmetry, instead of letting $\{a,b\}$ be the isolated 2-set, we can relabel variables and let any of $\{c,d\}$ or $\{e,f\}$ be the isolated 2-set and check for these properties.
-Hence, we can conclude that none of $a,b,c,d,e,f$ appears in any other set with each other (this is ruled out already for say $a,b$ by non-redundancy and remaining by negation of $\mathscr{P}(3,2)$ and symmetry).
-Also, we must have that each of $a,b,c,d,e,f$ must appear in at least one more 3-set.
-Since $\mathscr{P}(3,5)$ does not hold, it must be that both $a,b$ appear in exactly one other clause and by symmetry that each of $c,d,e,f$ appear in exactly one other clause.
-Also, since $\mathscr{P}(3,4)$ does not hold, we conclude that if the unique other 3-set where $a$ appears is say $\{a,x,y\}$ then it must be the case that there exists 3-set $\{c,x,y\}$.
-Symmetrically applying this argument to all pairs, we infer that case $\mathscr{P}(3,6)$ must hold for $F$, which is a contradiction.
-∎
-4.6.3.4 $s$ is even and either $a$ or $b$ appear in no other set  – $\mathscr{P}(3,1)$
-Without loss of generality let this be $a$. We take advantage of the fact that $a$ appears uniquely.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
-1 0 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
-0 1 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
-Total $\frac{20}{21}$
-4.6.3.5 $s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$  – $\mathscr{P}(3,2)$
-Without loss of generality let this be $a$ and $c$. Let this set be $\{a,c,x\}$.
-$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
+6.6.3.6 $\mathscr{P}(3\text{-}{ie},3)$$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in both of $cd$ and $ef$.
+Without loss of generality let this be $a$ and $c$, where $c=f$ and the additional two-clauses are $cd$, cece. Let this set be $\{a,c,x\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
-0 1 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$
-0 1 0 $d=x=1$ n5n-5 t3t-3 s4s-4 11 $\frac{5}{42}$
+0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$
+0 1 0 $d=e=x=1$ n6n-6 t4t-4 s6s-6 0 $\frac{1}{14}$
 Total 11
-4.6.3.6 $s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
+6.6.3.7 $\mathscr{P}(3\text{-}{ie},4)$$s$ is even, $|\{c,d,e,f\}|=3$ and there exists clause $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
 Without loss of generality let $v_{ab}=a$ and assume $c=f$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{4}{7}$
-0 1 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
 Total 11
-4.6.3.7 $s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,4)$
+6.6.3.8 $\mathscr{P}(3\text{-}{ie},5)$$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist clauses $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
 Without loss of generality let $v_{ab}=a$ and $v_{cdef}=c$.
-$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
@@
 Total 11
-4.6.3.8 $s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,5)$
+6.6.3.9 $\mathscr{P}(3\text{-}{ie},6)$$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist clauses $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
 Without loss of generality let $v_{ab}=a$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
 1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
-0 1 n2n-2 t1t-1 s1s-1 44 $\frac{17}{42}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 44 $\frac{17}{42}$
 Total 11
-4.6.3.9 $s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists  – $\mathscr{P}(3,6)$
-We directly compute:
-$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
-0 1 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
-1 0 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
-1 1 $x=y=0,c=d=e=f=1$ n8n-8 t6t-6 s10s-10 0 $\frac{1}{56}$
-Total $\frac{7}{8}$
 6.7 Proof of the Upper Bounds for $\Phi_{4}(s,t)$
-Let $F=F_{1}F_{2}$ for some monotone $3$-CNF $F_{2}$, where $F_{1}$ is composed of four distinct 2-clauses and one of them is $(ab)$.
-Let $\Xi$} be the class of such 3-set systems.
+Let $\mathbb{F}=(V,F_{1}\cap F_{2})\in{\Xi}_{4}(s,t)$ be a monotone $3$-CNF over $|V|$ variables with threshold $t$ and deficit $s$ where $F_{1}$ contains exactly four distinct 2-clauses and one of them is $\{a,b\}$.
 We cover $\Xi$} based on the following properties.
-(Note we only care about odd values of $s$, as even values are handled by the three 2-clause case.)
+Note we only care about odd values of $s$, as even values are handled by the three 2-clause case.
 $\mathscr{P}(4,1)$:
-No two two-clauses intersect.
+No two $2$-clauses intersect.
 $\mathscr{P}(4,2)$:
@@
 Some variable is in at least three 2-clauses.
-Lemma 4.13.
-Some $\mathscr{P}(4,j)$ must hold in a formula with at least have four two-clauses.
-Proof.
-Proceed by contradiction, assuming $\neg\bigwedge_{j=1}^{3}\mathscr{P}(4,j)$.
-In this case, by $\neg\mathscr{P}(4,1)$ some two-clause intersects another.
-This means some variable is in more than one 2-clause.
-By $\neg\mathscr{P}(4,2)$, it must be in at least three 2-clauses, implying $\mathscr{P}(4,3)$.
-∎
-4.7.0.1 No two two-clauses intersect  – $\mathscr{P}(4,1)$
-In this case, no three of the 2-clauses form a star, as they don’t overlap at all. We can therefore use the standard type 3 bound.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
+6.7.1 $\mathscr{P}(4,1)$ – No two $2$-clauses intersect
+Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,1)$. In this case, no three of the 2-clauses form a star, as they don’t overlap at all. We can therefore use the standard type 3 bound.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
 1 n1n-1 t1t-1 s2s-2 33 $\frac{9}{17}$
-0 1 n2n-2 t1t-1 s1s-1 33 $\frac{7}{17}$
+0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{7}{17}$
 Total $\frac{16}{17}$
-4.7.0.2 Some variable is in exactly two 2-clauses  – $\mathscr{P}(4,2)$
-Without loss of generality, assume $a$ is in both $ab$ and $ac$. Note $b$ and $c$ can be in as many 2-clauses as desired.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
+6.7.2 $\mathscr{P}(4,2)$ – Some variable is in exactly two 2-clauses
+Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,3)$. Without loss of generality, assume $a$ is in both $ab$ and $ac$. Note $b$ and $c$ can be in as many 2-clauses as desired.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
 1 n1n-1 t1t-1 s2s-2 22 $\frac{10}{17}$
@@
 Total $\frac{16}{17}$
-4.7.0.3 Some variable is in at least three 2-clauses  – $\mathscr{P}(4,3)$
-Without loss of generality, assume $a$ is in $ab$, acac, and $ad$. It may be in as many other 2-clauses as desired, as may $b$, $c$, and $d$.
-$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
+6.7.3 $\mathscr{P}(4,3)$ – Some variable is in at least three 2-clauses
+Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,3)$. Without loss of generality, assume $a$ is in $\{a,b\}$, $\{a,c\}$, and $\{a,d\}$.
+Core variables Forced variables Restricted CNF parameters Inductive bound
+$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
 1 n1n-1 t1t-1 s2s-2 0 $\frac{14}{17}$
@@
 Appendix A Optimality of Building Blocks
+Notation A.1.
+For a monotone CNF $\mathbb{F}=(V,F)$ and for $S\subset V$, we let $\mathbb{F}^{S}=(V-S,\mathbb{F}^{S})$ denote the CNF obtained by deleting the sets in $F$ that contain a variable in $S$, that is, $F^{S}=\{C\in F\mid C\cap S=\varnothing\}$.
+$\mathbb{F}^{S}$ corresponds to the CNF obtained by setting the variables in $S$ to 1.
+We use $\mathbb{F}^{\neg S}=(V-S,F^{\neg S})$ to denote the CNF obtained by deleting all occurrences of variables in $S$ from the clauses in $F$, that is, $F^{\neg S}=\{C-S\mid C\in F\}$.
+$\mathbb{F}^{\neg S}$ refers to the CNF obtained by setting the variables in $S$ to 0.
+For a variable $x$, let $\mathbb{F}^{x}=\mathbb{F}^{\{x\}}$ and $\mathbb{F}^{\neg x}=\mathbb{H}^{\neg\{x\}}$.
+For disjoint sets of variables $S$ and $T$, we use $\mathbb{F}^{S^{\prime},T^{\prime}}$ to denote be $S$ or $\neg S$ and $T^{\prime}$ can be $T$ or $\neg T$.
 A.0.1 ${\mathbb{T}}_{5}^{3}$ is $2$-extremal